### v1

# Web-CogReasoner: Towards Knowledge-Induced Cognitive Reasoning for Web Agents

Abstract
Multimodal large-scale models have significantly advanced the development of web agents, enabling them to perceive and interact with the digital environment in a manner analogous to human cognition. In this paper, we argue that web agents must first acquire sufficient knowledge to effectively engage in cognitive reasoning [^1] Drawing inspiration from Bloom’s educational philosophy, a cornerstone of modern pedagogy.. Therefore, we decompose a web agent’s capabilities into two essential stages: knowledge content learning and cognitive processes. To formalize this, we propose Web-CogKnowledge Framework, which categorizes knowledge into Factual, Conceptual, and Procedural domains. In this framework, knowledge content learning corresponds to the agent’s processes of Memorizing and Understanding, which rely on the former two types knowledge respectively, representing the "what" of learning. Conversely, cognitive processes correspond to Exploring, grounded in Procedural knowledge, defining the "how" of reasoning and action. To facilitate knowledge acquisition, we construct the Web-CogDataset, a structured resource curated from 14 real-world websites, designed to systematically instill the core knowledge necessary for web agent. This dataset serves as the agent’s conceptual grounding—the "nouns" upon which comprehension is built—as well as the basis for learning how to reason and act. Building on this foundation, we operationalize these processes through a novel knowledge-driven Chain-of-Thought (CoT) reasoning framework, developing and training our proposed agent, the Web-CogReasoner. Extensive experimentation reveals its significant superiority over existing models, particularly in its capacity for generalization to unseen tasks where its structured knowledge proves decisive. To facilitate rigorous and systematic evaluation, we introduce the Web-CogBench, a comprehensive evaluation suite designed to assess and compare agent performance across the delineated knowledge domains and cognitive capabilities. Our code and data is open sourced at https://github.com/Gnonymous/Web-CogReasoner.
footnotetext: These authors contributed equally to this work. Project Leader. Corresponding authors.
Figure 1: Visualization of the Web-CogKnowledge Framework along with the experimental results.
1 Introduction
The advent of large-scale models marks a landmark advancement in artificial intelligence. The development of Large Multimodal Models (LMMs), in particular, has vastly expanded the horizons for AI applications. In recent years, AI agents have emerged as a primary vehicle for deploying these models, significantly extending their capabilities. Much like humans, these agents can leverage a diverse array of tools to enhance their skills, leading to remarkable achievements in domains such as code generation hui2024qwen2 ; jiang2024survey , image huang2025diffusion ; bie2024renaissance and video synthesis  assran2025v ; liu2024sora , and academic research. Notably, the application of web agents to surfing the internet has also seen substantial progress.
The evolution of web agents has transitioned from early rule-based systems to modern approaches that widely leverage Large Language Models (LLMs) or Languge Vision Models (LVMs) wang2024survey ; ning2025survey ; zhang2024large ; sapkota2025ai . These LLM-powered agents typically process HTML or Accessibility Tree inputs, formulating tasks and observations into natural language prompts for the model to reason upon and decide the next action. With the integration of LVMs, web agents have gained perceptual abilities akin to human vision, enabling them to comprehend and reason about multimodal content on web pages. As a whole, web agents can be broadly categorized into three types: (1) Text-only agents zhou2023webarena ; li2023camel , which suffer from information loss due to their inability to capture visual cues; (2) Vision-only agents qin2025ui , which rely solely on visual information and thus lose critical structured data; and (3) Hybrid agents koh2024visualwebarena ; he2024openwebvoyager , which utilize both visual and textual information.
LLMs and LMMs are pre-trained on vast amounts of general-domain knowledge, which is undeniably crucial for their foundational capabilities. However, as with any specialized field, this general knowledge is often insufficient, creating a performance bottleneck that caps the potential of the three aforementioned model types. While many existing works have attempted knowledge enhancement or injection, these methods often lack a systematic or theoretically grounded approach. To address this, we draw inspiration from a highly successful paradigm in human education: Bloom’s Taxonomy ormell1974bloom ; conklin2005taxonomy . This framework suggests that effective learning can be deconstructed into two key stages: knowledge content learning and cognitive processes. Accordingly, we propose that training a web agent should follow a similar two-stage methodology. The first stage is dedicated to acquiring knowledge of varying difficulties. The second stage then focuses on the analysis, application, and ultimately, the creative synthesis of that knowledge. This mirrors the human learning trajectory: we first accumulate knowledge through education (Stage 1), and then based on that foundation of knowledge and experience, we learn to apply, innovate, and create (Stage 2).
Therefore, in this paper, we introduce a training paradigm for our model grounded in Bloom’s Taxonomy, bifurcated into two sequential stages: Knowledge Content Learning and Cognitive Processes. The initial stage, Knowledge Content Learning, is dedicated to building a multi-layered knowledge foundation. This foundation consists of three distinct types: Factual knowledge, which includes basic, foundational concepts; Conceptual knowledge, which maps the interrelationships between these facts; and Procedural knowledge, which represents the logical frameworks needed to synthesize factual and conceptual understanding for effective reasoning. To curate this knowledge, we construct Web-CogDataset, collecting metadata from 14 prominent websites and engineered 12 granular tasks designed to systematically instill each knowledge type. In essence, this stage teaches the agent foundational web knowledge and how to apply it to familiar tasks. The subsequent stage, Cognitive Processes, focuses on cultivating the agent’s self-exploring capability. This advanced capability empowers the agent to build upon its learned foundation, enabling it to "learn how to learn" by creatively exploring its own "self-knowledge" to tackle previously unencountered tasks and challenges. To operationalize this, we developed knowledge-guided reasoning templates and utilized imitation learning to train the model. This process explicitly teaches the agent the cognitive faculties required for autonomous learning and creative problem-solving in novel scenarios. To validate our approach, we conducted rigorous evaluations on a suite of general-purpose benchmarks as well as our own. Our findings reveal that the agent trained with our methodology consistently and significantly outperforms state-of-the-art baselines. Crucially, and in alignment with our hypothesis, the performance gap widens dramatically on knowledge-intensive tasks. This demonstrates that, analogous to human learning, the structured acquisition of knowledge is a key determinant of an agent’s ability to excel in complex, domain-specific scenarios.
In summary, our contributions are threefold:
  1. 1.
    Drawing inspiration from Bloom’s taxonomy and established human educational paradigms, we propose the Web-CogKnowledge Framework — a systematic, two-stage training methodology designed to enhance the cognitive capabilities of web agents. Built upon this framework, we develop Web-CogReasoner. Rigorous benchmarking demonstrates that agents trained under our framework achieve a significant performance improvement over current state-of-the-art models.
  2. 2.
    We construct the Web-CogDataset, a structured curriculum consisting of 12 fine-grained and progressively challenging tasks. These tasks are meticulously designed to incrementally build the agent’s web knowledge, cognition capability, and higher-order reasoning.
  3. 3.
    To enable comprehensive and robust evaluation, we introduce Web-CogBench, a novel benchmark specifically designed to assess whether a web agent possesses the requisite prior knowledge and cognitive capabilities for effective web navigation. This benchmark will be released publicly to foster further research in this area.
2 Related Work
Web Agent
Early studies on web page understanding have concentrated on leveraging structured text. The work gur2022understanding defined crucial HTML understanding tasks, such as semantic classification, description generation, and autonomous web navigation, and trained models to comprehend the structure and content of HTML. AutoWebGLMlai2024autowebglm introduced an HTML simplification algorithm. Through curriculum learning, it first trained the model using collected web page data to recognize web page structures and understand the functions of different interactive components, like text boxes and buttons. Subsequently, the model learned to execute simple web tasks and decompose complex ones into subtasks, enabling the complete execution of web tasks from simple to complex vision-based Web Page Understanding.
Recent research has explored the use of visual information for web page understanding. SeeClickcheng2024seeclick designed specialized pre-training tasks, including text-to-point and text-to-bounding box localization tasks, as well as point-to-text and bbox-to-text generation tasks. These tasks facilitated the model’s learning of the correlation between the positions of web elements and their textual descriptions, enhancing the model’s ability to localize and understand web elements. CogAgenthong2024cogagent introduced a high-resolution cross-module, enabling the model to recognize tiny elements and text in GUIs, and constructed a large-scale GUI annotated dataset for GUI grounding tasks, achieving excellent performance in multiple visual question answering (VQA) benchmarks and GUI navigation task datasets. OmniParserwan2024omniparser integrated three typical visually-situated text parsing tasks—text spotting, key information extraction, and table recognition—within a unified framework, providing a more efficient solution for GUI Agents to handle complex tasks. Unlike many existing agent frameworks that rely on heavily wrapped commercial models (e.g., GPT-4o) with expert-crafted prompts and workflows, UI-TARSqin2025ui adopted an end-to-end architectural design, directly perceiving screenshots, executing reasoning processes, and autonomously generating valid actions. UGroundgou2024navigating developed a universal visual grounding model using a GUI dataset with 10 million elements, demonstrating outstanding performance in desktop tasks and mobile tasks.
Multimodal approaches have emerged as a powerful way to comprehensively understand web pages. WebVoyagerhe2024webvoyager utilized the GPT-4V-ACT tool to overlay bounding boxes and numerical labels of interactive elements on web page screenshots. Additionally, it complemented web page visual information with the Accessibility Tree, helping the model gain a more comprehensive understanding of web page content. SeeActzheng2024gpt directly harnessed the powerful visual and language understanding capabilities of GPT-4V and explored various grounding strategies to map the text plans generated by the model into specific web operations. Tuning-free Attention-driven Grounding (TAG) xu2025attention method further accomplishes this task without the need for additional fine-tuning by leveraging the inherent attention patterns in pretrained MLLMs.
Web Agent Evaluation
To evaluate GUI agents’capabilities in web interactions, benchmarks are broadly categorized into web browsing tasks and web understanding assessments. For web browsing, offline benchmarks like Mind2Webdeng2023mind2web , Multimodal-Mind2Web, AutoWebBenchlai2024autowebglm and WebVLN-v1chen2024webvln evaluate models’ability to perform multi-step web tasks (e.g., booking flights, filling dynamic forms) using static datasets that include webpage content along with corresponding interaction trajectories. In contrast to offline benchmarks, online environments such as Mini-WoB++liu2018reinforcement , Webshopyao2022webshop , and WebArenazhou2023webarena enable real-time interaction testing. Mini-WoB++liu2018reinforcement focuses on low-level UI operations (e.g., clicking buttons, handling pop-ups) within simplified webpages. By comparison, Webshopyao2022webshop and WebArenazhou2023webarena replicate complex, real-world tasks—such as e-commerce transactions and Wikipedia editing—to evaluate task completion performance. Building on this, VisualWebArenakoh2024visualwebarena incorporates multimodal inputs to assess models’ ability to handle vision-augmented interactions in dynamic web environments. To evaluate deeper web understanding, benchmarks are designed to target specific cognitive abilities. WEBQAchang2022webqa assesses a model’s capacity for open-domain, multi-hop reasoning. In contrast, ScreenQAhsiao2022screenqa and the ScreenAIbaechler2024screenai suite focus on screen-centric comprehension. ScreenQAhsiao2022screenqa evaluates a model’s ability to recognize UI elements and interpret context by answering questions based on screenshots. ScreenAIhsiao2022screenqa further extends this with three progressively challenging subtasks: Screen Annotation (SA), which involves spatial layout tagging; ScreenQA Short, which tests concise fact extraction; and Complex ScreenQA, which requires multi-step reasoning across diverse screen layouts. Together, these tasks assess fine-grained layout understanding, semantic interpretation, and practical problem-solving in visually dense interfaces.
3 Web-CogKnowledge Framework
3.1 Bloom’s Taxonomy
The Bloom’s Taxonomy  [^2] https://fctl.ucf.edu/teaching-resources/course-design/blooms-taxonomy/. (Anderson & Krathwohl, 2001) presents a two-dimensional pedagogical design, i.e., knowledge content learning and cognitive process dimension, which demonstrates a "shallow-to-deep" instructional methodology. This approach posits that effective learning is not monolithic but is a structured progression through distinct types of knowledge, beginning with foundational elements and culminating in high-level self-awareness. This educational strategy ensures that learners first establish a solid base of facts and concepts before advancing to the execution of complex procedures and, ultimately, the application of reflective metacognitive strategies.
This pedagogical progression from simple to complex can be formally outlined through four hierarchical types of knowledge: Factual Knowledge, Conceptual Knowledge, Procedural Knowledge and Metacognitive Knowledge. More details can be found in Appendix A.1.1.
3.2 Web-CogKnowledge
Motivated by Bloom’s taxonomy, we propose a hierarchical web knowledge framework that structures web knowledge according to its taxonomy, collects corresponding knowledge at each level, and trains the model correspondingly. We refer to this knowledge as Web-CogKnowledge. This Web-CogKnowledg decomposes knowledge into three levels:
  1. 1.
    Factual knowledge: concrete information extracted from web content. Such as identifying the attributes of individual web elements and predicting the immediate, direct consequences of a single interaction;
  2. 2.
    Conceptual knowledge: semantic relationships and abstract patterns underlying webpage content and structure. Such as inferring the function of interface components, comprehending the overall purpose and structure of a webpage, and interpreting its multimodal content;
  3. 3.
    Procedural knowledge: actionable know-how for accomplishing specific tasks through interaction, including planning, decision-making, and sequential execution. Such as executing goal-oriented action sequences, inferring user intent from observed behaviors, and handling unexpected interruptions to complete complex tasks;
This taxonomy aligns each knowledge type with a corresponding cognitive competency required for web-based reasoning and interaction.
Figure 2: Statistics of Selected Websites by Category.
To fully leverage the potential of Web-CogKnowledge, leading to the ultimate realization of the Web-CogKnowledge Framework. We first collect multimodal metadata from 14 real-world websites (see Figure 2) and then design Web-CogDataset with diverse set of web tasks for Web-Agent training in Section 3.3. And finally construct Web-CogBench in Section 3.4, built from meticulously chosen subsets. Specifically, we curate the metacognitive knowledge and Web-CogReasoner learning and training process in Section 4 due to its importance and complexity. More details about data collection and cleaning process can be found in Appendix A.1.3.
3.3 Web-CogDataset
By selectively crawling metadata from the aforementioned websites and leveraging the hierarchical structure of Web-CogKnowledge, we define a series of tasks corresponding to different levels of knowledge granularity. This process results in the construction of a comprehensive and hierarchically structured dataset, referred to as Web-CogDataset. The task definitions, organized according to the knowledge hierarchy, are introduced below. Detailed statistics and illustrative examples for Web-Dataset are provided in Table 1 and Figure 3. For additional implementation details, please refer to the Appendix A.2.
Table 1: Statistics of the Web-CogDataset.
Knowledge Task Statistics
Subtotal Total
Factual Web Knowledge Element Attribute Recognition 37K 81K
Sub-elements Prediction 1K
Page Change Prediction 24K
Next Page Prediction 18K
Source Element Prediction 1K
Conceptual Web Knowledge Element Understanding 40K 62K
WebPage Understanding 7K
Caption & QA 15K
Procedural Web Knowledge User’s Intention Prediction 2K 27K
Popup Close 1K
Single-Step Web Task 17K
Noisy Multi-Steps Web Task 7K
Factual Web Knowledge Tasks Overview
The Factual Web Knowledge tasks are designed to establish the agent’s core perceptual abilities by operationalizing our definition of factual knowledge. These five tasks train the agent to identify the attributes of individual web elements and predict the immediate, direct consequences of a single interaction. By focusing on extracting concrete, visually grounded information—rather than abstract reasoning—this task category ensures the agent builds a foundational understanding of the web’s basic cause-and-effect mechanics. This layer of knowledge is the essential prerequisite for any subsequent conceptual or procedural reasoning.
  • Element Attribute Recognition: Given a screenshot with a highlighted interactive element, the model predicts its semantic role (e.g., button, link) and accessible name (e.g., "Submit", "Search"), relying solely on visual cues.
  • Sub-elements Prediction: The model infers the dynamic hierarchical structure of a web interface by predicting the set of sub-elements that appear upon interacting with a highlighted parent element.
  • Page Change Prediction: The model anticipates the visual changes that occur after clicking on a specific element, formulating an open-ended response that describes the expected dynamic behavior of the page.
  • Next Page Prediction: The model predicts the subsequent page that results from interacting with a specific element on the current page. To enhance generalization, we designed two types of tasks: multiple-choice questions and open-ended responses.
  • Source Element Prediction: Given two screenshots—the current page and the resulting target page. Then the model identifies which of the visually marked elements on the current page leads to the target, simulating visual cause-and-effect reasoning.
Figure 3: Illustration of Web-CogDataset.
Conceptual Web Knowledge Tasks Overview
Building upon the factual foundation, the Conceptual Web Knowledge tasks are designed to move the agent from perception to comprehension. These tasks train the agent to synthesize discrete facts into meaningful patterns and understand the semantic relationships underlying a webpage’s content and structure. Rather than merely identifying elements, the agent learns to infer their functions, comprehend the webpage’s overall purpose and organization, and interpret its multimodal content. This layer of conceptual understanding is critical, as it enables the agent to grasp the "why" behind the interface, transforming it from a collection of static elements into a coherent, interactive system.
  • Element Understanding: For a specific interactive element, the model generates an open-ended paragraph that comprehensively describes the element’s Visible Traits (e.g., text, shape, styling), its On-page Location (e.g., header, sidebar, main content), and its likely User-facing Function (e.g., playing a video, navigating to a new page), relying solely on visual context.
  • WebPage Understanding: Given a full-page screenshot, the model generates a comprehensive overview describing the webpage’s Layout Organization (e.g., header, event information, seating chart, filter panel), Key Element Analysis (e.g., element attribute, description, function, Interaction, expected outcome), and a Summary of the WebPage. This enables a thorough understanding of the webpage’s structure and functionality.
  • Caption & QA: The model generates captions and answers questions about embedded images and WebPages, assessing its visual reasoning and holistic understanding of multimodal webpage information (e.g., layout, text). Contributes to further enhancing the model’s understanding of WebPages, while preserving its open conversational abilities with web content. The orignal data is drawn from the Multi-UI liu2024harnessing dataset.
Procedural Web Knowledge Tasks Overview
With both factual and conceptual knowledge established, the Procedural Web Knowledge tasks are designed to transition the agent from understanding to action. These tasks train the agent to formulate and execute goal-oriented plans by leveraging its actionable know-how for accomplishing specific tasks through interaction. The agent learns to decompose complex user goals into sequential steps, handle unexpected interruptions like popups, and orchestrate a series of interactions to achieve a desired outcome. This final knowledge layer is paramount, as it empowers the agent to apply its understanding strategically, converting passive knowledge into active, successful task completion.
  • User’s Intention Prediction: The model infers high-level user intent from a sequence of webpage screenshots representing an interaction trajectory, requiring visual understanding and temporal reasoning. The task is built on the MultiModal-Mind2Web deng2023mind2web dataset, mapping screenshot sequences to natural language instructions.
  • Popup Close: The model identifies and dismisses popups (e.g., notification modals, login forms) on synthesized webpage screenshots, using a dataset of 51 popup components from JS Design  [^3] https://js.design/. overlaid on OpenWebVoyager he2024openwebvoyager webpages, with combinatorial augmentation of closing strategies for diverse training.
  • Single-Step Web Task: The model executes goal-directed web interactions to perform atomic web actions, using annotated screenshots from the MultiUI liu2024harnessing dataset, with tasks validated for accuracy.
  • Noisy Multi-Steps Web Task: The model perform complex, goal-oriented interactions across multiple steps, based on unified OpenWebVoyager he2024openwebvoyager trajectories, with a standardized action space for consistent evaluation. To test robustness, some tasks include UI interruptions (e.g., popups) that must be dismissed before resuming the task.
Table 2: Detailed statistics of Web-CogBench.
Task Congition Metric #Num
Element Attribute Recognition Memorizing ROUGE-L 249
Next Page Prediction Accuracy 93
Source Element Prediction Accuracy 32
Element Understanding Understanding LVM Judge 200
WebPage Understanding LVM Judge 77
User’s Intention Prediction Exploring LVM Judge 105
Popup Close Accuracy 58
Single Step Exploration Accuracy 62
Total - - 876
3.4 Web-CogBench
To rigorously evaluate the cognitive capabilities fostered by our knowledge-centric framework, we introduce Web-CogBench. While our training dataset is structured by knowledge types (Factual, Conceptual, Procedural), Web-CogBench assesses agent performance through the lens of three corresponding cognitive abilities: Memorizing, Understanding, and Exploring. This benchmark is curated from a representative subset of our Web-CogDataset and is designed to measure how well an agent translates its learned knowledge into practical skills within complex web contexts. Detailed statistics of Web-CogBench are provided in Table 2. The three evaluation dimensions directly mirror our hierarchical knowledge framework:
Memorizing
Assessing the agent’s ability to recall and recognize concrete information, directly corresponding to the acquisition of Factual Knowledge. It evaluates whether the agent can accurately identify the attributes of web elements and the state of a webpage.
Understanding
Measuring the agent’s capacity for semantic interpretation, aligning with the mastery of Conceptual Knowledge. It tests whether the agent can move beyond mere identification to comprehend the function of elements and the contextual relationships within a page.
Exploring
Evaluating the agent’s ability to plan and execute goal-oriented actions, reflecting the application of Procedural Knowledge. It assesses whether the agent can strategically navigate the web, handle interruptions, and complete multi-step tasks to fulfill user goals.
4 Web-CogReasoner
4.1 Problem Setup
We model the interaction between the environment and Web-CogReasoner as a partially observable Markov decision process (POMDP): $P=(S,A,O,K,T,R)$, where $S$ denotes the webpage status, $A$ is the action space (See complete operations in Table 1). $O$ is the observation space, and $K$ represents the internal knowledge of Web-CogReasoner. $T:S\times A\rightarrow S$ is the transition function.
[…]
$(h_{t},a_{t})=\pi_{\theta}(\cdot|K,I,Q,o_{1},h_{1},a_{1},\dots,o_{t})$
Finally, the overall performance is assessed using a binary reward function $R$, which returns 1 only if the agent successfully completes the user-specified task at the final step, and 0 otherwise.
4.2 Web-CogReasoner Framework
We introduce Web-CogReasoner, a knowledge-driven reasoning system built on LVMs and trained on the Web-CogDataset in Section3.3. It solves complex web tasks by performing chain-of-thought reasoning grounded in knowledge acquired during training. We define web agent actions set based on common user behaviors observed during web browsing. At each interaction step, the model receives a screenshot of the current webpage along with its corresponding AX Tree. The model integrates visual, structural representations with other contextual cues and history trajectories to predict the next action. Once an action is predicted, it is executed by an automated web agent. This process iterates until the task is completed or reach the maximum step limit. The following section details the design of our agent and reasoning framework:
Observation Space
To effectively understand and interact with complex web interfaces, our web agent takes as input both webpage screenshots and corresponding AX Tree. The screenshot provides rich visual context essential for grounding visual reasoning, while the AX Tree offers a structured semantic representation of page elements. Specifically, the AX Tree encodes key properties such as role (indicating the type of an element), name (how the element should be referred to), description (additional descriptive content), and state (e.g., whether a checkbox is checked or a dropdown is expanded).
Action Space
To enable web agents to operate in real-world environments, we define an action space (Table 3) grounded in common user interactions during web browsing. Each action $a_{t}$ operates on a target element identified by a unique ID from the current observation $x_{t}$. For agents using AX Tree representations, this ID refers to a node within the tree.
Table 3: Action Space of Web-CogReasoner for Web Interaction.
Instruction Description
click [id] Click an element
type [id] [content] Input specified content into an element
scroll [id or WINDOW] [up/down] Scroll an element or the page up/down
dbclick [id] Double-click an element
go_back Navigate to the previous webpage
go_forward Navigate to the next webpage
stop [content] Submit the final answer
Restart Restart the current task
Wait Wait for one second before proceeding
Knowledge-driven CoT Reasoning
Figure 4: An overview of the Web-CogReasoner framework, illustrating the Knowledge-driven Chain of Thought (CoT) process. The agent initiates with a Task and observes the webpage (Observe Space). The Knowledge-driven COT Reasoning module then systematically leverages different types of Web-CogKnowledge—Factual Knowledge (blue), Conceptual Knowledge (green), and Procedural Knowledge (yellow). This structured reasoning guides the Planning phase, which decomposes the task and formulates a step-by-step strategy. The process culminates in a concrete Action to be executed on the webpage.
To instantiate the hierarchical Web-CogKnowledge Framework in a reasoning-capable agent, we introduce the Web-CogReasoner, a structured cognitive web-agent that generates a Knowledge-driven Chain-of-Thought (CoT) to guide decision-making in web environments. As illustrated in Figure 4, this CoT is not a monolithic or opaque text block; rather, it is a scaffolded reasoning template, where each segment is explicitly grounded in a distinct type of web knowledge. This design ensures interpretability, traceability, and faithful knowledge alignment throughout the reasoning process.
The CoT construction follows a layered progression, with each stage corresponding to a cognitive function defined in the Web-CogKnowledge taxonomy:
  • Factual Knowledge (highlighted in blue) constitutes the foundation of reasoning by anchoring the model in observable page facts. This phase populates the Webpage Layout Description and Key Element Analysis sections. The agent extracts concrete interface features such as DOM structure, ARIA attributes (e.g., role=‘search’), accessible names, and current page states. At this stage, the agent is essentially answering the question: “What is on the page?”
  • Conceptual Knowledge (highlighted in green) builds upon the factual layer by introducing semantic understanding. Within the same sections, the model interprets the purpose of key UI components and predicts their behavior under user interaction. For instance, it explains the functional intent of the search bar (“allows users to input a location”) and anticipates changes to the page layout (“the list of stores will update after input”). This phase reflects the agent’s capacity to comprehend: “What does this element mean?” and “How does it contribute to the task?”
  • Procedural Knowledge (highlighted in yellow) informs goal-oriented planning and task decomposition. It is primarily instantiated in the Task Recap, Task Decomposition, Step-by-Step Reasoning, and Final Action Summary sections. The Reasoner infers the user’s intent from the task prompt, predicts the next sub-goal for interaction (e.g., “Users will type in their desired location”), and generates a sequential plan:
    1. 1.
      Enter the ZIP code 90028 into the search bar;
    2. 2.
      Review the updated list of Apple Stores;
    3. 3.
      (Fallback) Click the “Complete store list” if results are missing.
    This layer translates semantic understanding into executable strategies—effectively answering “How should the task be accomplished?”
By modularizing the CoT along these knowledge axes, Web-CogReasoner achieves structured planning grounded in a cognitively coherent framework. Unlike conventional end-to-end models that produce opaque or hallucinated justifications, each reasoning step here is explicitly traceable to factual evidence, semantic inference, or procedural logic. This alignment reduces ungrounded output and enhances interpretability.
As shown in Figure 4, the color-coded CoT segments (yellow, purple, blue) illustrate how the agent transitions from perception to comprehension to action. This workflow encapsulates the chain:
[…]
Through this process, the Web-CogReasoner becomes a practical embodiment of Web-CogKnowledge, leveraging its hierarchical structure to support robust, transparent, and goal-directed web reasoning.
4.3 Training
We employ a multi-stage Imitation Learning strategy to train our model on the Web-CogDataset, utilizing Qwen2.5-VL-7B bai2025qwen2 as the base model. Each stage is aligned with a distinct layer of the Web-CogKnowledge Framework: (1) the first knowledge content learning focuses on acquiring Factual Knowledge and Conceptual Knowledge, enabling the model to interpret web content and semantics; (2) the second cognitive process emphasizes Procedural Knowledge, training the model to plan and execute multi-step web interactions. To accommodate the increased complexity of the final stage, which involves multi-image inputs and extended reasoning, we configure training with a maximum sequence length of 8K and a batch size of 1 with gradient accumulation of 16 steps. All experiments are conducted on a cluster of 8 x NVIDIA A800 80GB GPUs.
5 Experiment
5.1 Experimental Setup
5.1.1 Models and Baselines
To comprehensively evaluate Web-CogReasoner, we compare its performance against several representative and state-of-the-art baselines.
Our proposed model, Web-CogReasoner, is built upon the Qwen2.5-VL-7Bbai2025qwen2 large multimodal model (LMM) as its foundational architecture. As detailed in Section 4, Web-CogReasoner extends Qwen2.5-VL by integrating a knowledge-induced cognitive reasoning framework, which leverages a structured knowledge base (described in Section 3) to guide chain-of-thought (CoT) reasoning for more robust and generalizable web navigation. Specifically, we performed supervised fine-tuning on Qwen2.5-VL using our curated datasets to imbue it with the proposed knowledge-driven capabilities.
[…]
These baselines collectively represent a diverse spectrum of current web agent methodologies, including foundational LLMs, interactive LLM-based agents, and powerful general-purpose LLM/LMMs leveraging prompting strategies. This allows for a thorough and multifaceted evaluation of Web-CogReasoner’s effectiveness and its unique advantages conferred by knowledge-induced cognitive reasoning.
5.1.2 Datasets
Training Dataset
In the training phase, we utilized a comprehensive cognitive dataset——Web-CogDataset,which is meticulously constructed as detailed in Section 3.3. This dataset is specifically designed to imbue Web-CogReasoner with knowledge-induced cognitive reasoning capabilities. It encompasses a rich variety of samples aligned with Factual, Conceptual, and Procedural knowledge dimensions, including diverse web page screenshots, corresponding accessibility trees, and detailed interaction trajectories with associated reasoning chains. Its comprehensive and structured nature is crucial for enabling the model to learn and perform sophisticated web navigation tasks.
Evaluation Datasets
To rigorously evaluate Web-CogReasoner’s performance and generalization capabilities, we employ a multi-faceted evaluation strategy utilizing various prominent datasets. We utilize the following datasets for evaluation:
  • Web-CogBench: As elaborated in Section 3.4, this custom-built benchmark is meticulously designed to assess distinct knowledge-driven cognitive abilities: Memorizing, Understanding, and Executing. Aallowing for granular evaluation of how our knowledge-induced reasoning framework contributes to these specific competencies.
  • VisualWebBench liu2024visualwebbench : A human-curated multimodal benchmark covering 1.5K examples across 139 real websites and 87 sub-domains, designed to evaluate MLLMs’s grounding and reasoning abilities in complex, web-based scenarios.
  • WebVoyager he2024openwebvoyager : This dataset comprises tasks on 15 websites that were seen during training and includes a total of 643 online web task queries. It primarily assesses the agent’s ability to perform tasks on familiar web environments.
  • Online Multimodal-Mind2Web deng2023mind2web : This includes two complementary subsets. The Cross-Task Test Set contains 112 queries from 33 websites seen during training, evaluating generalization to new tasks within known structures. The Cross-Website Test Set comprises 106 queries from 6 unseen websites across familiar domains (Entertainment, Shopping, and Travel), assessing the agent’s ability to transfer domain knowledge to novel layouts.
5.2 Evaluation Metrics
To provide a comprehensive and quantitative assessment of our model’s performance, we employ a set of specific metrics tailored to the nature of our evaluation datasets.
Web-CogBench Metrics
For our Web-CogBench, which evaluates distinct cognitive abilities, we use a variety of metrics depending on the task format:
  • Accuracy: For tasks with a definitive correct answer, such as multiple-choice questions (Next Page Prediction) or single-element identification (Source Element Prediction, Popup Close, Single Step Exploration), we calculate the standard accuracy. An answer is considered correct only if it exactly matches the ground truth.
  • ROUGE-L: For generative tasks that produce open-ended text, such as Element Attribute Recognition, we use ROUGE-L to measure the F1-score of the longest common subsequence between the predicted text and the ground truth. This metric effectively evaluates content overlap for short-form text generation.
  • LVM Judge: For tasks requiring qualitative assessment of generated content, such as Element Understanding and WebPage Understanding, a simple string match is insufficient. Therefore, we employ a powerful Large Vision Model (LVM) as an automated judge. The LVM is prompted to evaluate the quality, relevance, and accuracy of the model’s generated description against the ground truth, returning a numerical score on a 5-point scale. This approach provides a scalable and consistent method for evaluating complex, open-ended visual reasoning.
Online Web Tasks Metric
For online evaluation on benchmarks like WebVoyager and Multimodal-Mind2Web, the primary metric is the Task Success Rate (SR). A task is considered a success if and only if the agent successfully completes all required sub-steps and achieves the final goal specified in the instruction within the maximum allowed number of steps. The final success rate is the percentage of successfully completed tasks over the total number of tasks in the test set. This metric provides a clear and direct measure of the agent’s practical utility and effectiveness in real-world, dynamic environments.
5.3 Main Result
To validate the effectiveness of our proposed Web-CogReasoner and its Web-CogKnowledge Framework inspired by Bloom’s taxonomy, we conduct a comprehensive set of experiments across several challenging benchmarks. We evaluate our model’s performance in three key areas: fine-grained cognitive capabilities, visual understanding, and practical generalization on live online tasks. Our model is compared against leading proprietary agents, such as Claude Sonnet 4 and Gemini 2.5 Pro, as well as state-of-the-art open-source models.
Table 4: Performance evaluation on the Web-CogBench benchmark.
[…]
Overall
Claude Sonnet 4 54.3 64.7 100 96.8 76.8
Gemini 2.5 Pro 73.5 51.9 96.6 98.4 80.2
Qwen2.5-VL-7B 62.0 51.9 91.4 90.3 69.8
UI-TARs-7B-SFT 48.0 32.4 25.9 33.9 46.4
Web-CogReasoner (Ours) 79.0 61.4 98.3 95.2 84.4
Results on Web-CogBench
We first evaluate Web-CogReasoner on our proposed Web-CogBench, a benchmark specifically curated to assess the foundational knowledge and cognitive reasoning capabilities of web agents. It encompasses twelve diverse tasks ranging from element recognition to user intent inference, in alignment with our two-stage cognitive framework.
As presented in Table 4, Web-CogReasoner achieves state-of-the-art performance with an overall accuracy of 82.9%, surpassing both commercial and open-source baselines. Notably, it outperforms the powerful proprietary model Gemini 2.5 Pro (77.6%) and the strong open-source model Qwen2.5-VL-7B (66.9%), with substantial gains in high-level reasoning tasks such as Source Element Prediction (87.5% vs. 84.4%) and WebPage Understanding (79.0% vs. 73.5%). These results confirm the effectiveness of our design, which couples structured knowledge acquisition (factual, conceptual, procedural) with Knowledge-driven Chain-of-Thought (CoT) reasoning to form a solid cognitive foundation for agents.
Importantly, we find that strong performance on Web-CogBench is highly indicative of overall agent effectiveness across web tasks. For example, Gemini performs best on Web-CogBench among baseline models and also exhibits the best overall task completion rate in live-agent evaluations, in table 6 and table 7. However, we further observe that foundational model capabilities—particularly in visual perception—play a critical role in enabling such high-level cognition. Agents with weaker visual recognition skills exhibit reduced performance even when endowed with strong reasoning mechanisms, as reflected in additional evaluation results from VisualWebBench, in table 5.3, especially in perception-oriented tasks.
In summary, Web-CogBench serves not only as a direct measure of an agent’s reasoning capacity, but also implicitly requires strong perceptual grounding. Excelling on Web-CogBench implies that an agent possesses both the knowledge reasoning ability and the necessary perceptual competence—making it a highly reliable proxy for real-world web exploration effectiveness.
Table 5: Performance evaluation on the VisualWebBench benchmark. Tasks are grouped into Perception-Oriented and Reasoning-Oriented categories.
[…]
Ground
Claude Sonnet 4 73.3 72.6 96.2 80.7 81.1 96.1 96.3 91.2 85.9
Gemini 2.5 Pro 74.9 70.8 95.1 80.3 91.8 96.8 90.3 92.9 86.6
Qwen2.5-VL-7B 70.8 71.7 81.4 74.6 77.5 86.8 68.0 77.4 76.0
UI-TARs-7B-SFT 71.3 78.7 97.2 82.4 91.8 91.8 85.4 89.7 86.0
Web-CogReasoner (Ours) 67.2 72.6 97.0 79.0 96.4 96.1 88.4 93.6 86.3
Results on VisualWebBench
To evaluate the model’s visual understanding capabilities, we conduct experiments on the VisualWebBench benchmark liu2024visualwebbench . The results, summarized in Table 5.3, expose the limitations of visually-focused benchmarks and motivate the development of our proposed Web-CogBench.
Our Web-CogReasoner achieves the highest average score of 86.3%, marginally outperforming UI-TARs (86.0%). While this may suggest comparable performance at first glance, a closer examination reveals a stark contrast: UI-TARs performs significantly worse on Web-CogBench, with an overall score of only 48.2% (see Table 4).
This discrepancy underscores a critical insight — excelling in tasks such as OCR and element localization does not imply competence in cognitive reasoning. High scores on visual benchmarks like VisualWebBench can obscure deeper limitations in understanding and decision-making. In contrast, Web-CogReasoner demonstrates consistent strength across both visual and reasoning benchmarks, confirming that our framework effectively integrates visual perception with advanced cognitive reasoning. This dual capability is essential for building robust and reliable web agents.
Results on Online Web Tasks
Beyond offline benchmarks, we evaluated Web-CogReasoner’s practical utility and generalization ability on live, online web tasks using the WebVoyager he2024openwebvoyager and Online Multimodal-Mind2Web deng2023mind2web datasets. These experiments test the agent’s ability to apply its knowledge and reasoning to complete complex, multi-step and live tasks on unseen websites and scenarios, which is the ultimate goal for any autonomous web agent. Notably, we did not include UI-TARs-7B-SFT qin2025ui in the online evaluation due to the absence of publicly available online inference scripts. Therefore, we focus our comparison on the OpenWebVoyager he2024openwebvoyager series, which are specifically designed for web task execution.
Table 6: Task success rates on the WebVoyager test set. We compare our model, Web-CogReasoner, against several baselines across 14 websites. The "Overall" score is the average success rate.
Agent Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
Claude Sonnet 4 26.7% 87.8% 48.8% 69.8% 68.3% 2.3% 45.5% 83.3%
Gemini 2.5 Pro 60.0% 63.4% 62.8% 67.4% 68.3% 9.1% 56.8% 73.8%
Qwen2.5-VL-7B 0.0% 0.0% 0.0% 4.7% 0.0% 0.0% 0.0% 2.3%
[…]
Web-CogReasoner (Ours) 26.7% 31.7% 32.6% 34.9% 29.3% 2.3% 15.9% 54.8%
[…]
Overall Alpha
Claude Sonnet 4 23.8% 37.2% 4.8% 80.5% 48.8% 82.6% 47.7%
Gemini 2.5 Pro 52.3% 76.7% 4.8% 75.6% 58.1% 82.6% 54.9%
Qwen2.5-VL-7B 0.0% 11.6% 0.0% 2.4% 7.0% 2.2% 2.2%
[…]
Web-CogReasoner (Ours) 14.3% 55.8% 9.5% 39.0% 37.2% 39.1% 30.2%
[…]
Table 7: Performance comparison on the Mind2Web Online dataset under cross-task and cross-websites
Agent Cross-task (Unseen Task) Cross-web (Unseen Websites)
Entertainment Shopping Travel Overall Entertainment Shopping Travel Overall
Claude Sonnet 4 44.9% 35.3% 40.0% 40.2% 45.5% 6.7% 14.0% 21.7%
Gemini 2.5 Pro 46.9% 35.3% 28.3% 37.5% 42.4% 10.0% 23.3% 25.5%
OpenWebVoyagerMax 22.4% 29.4% 15.2% 20.5% 3.0% 8.7% 23.3% 11.7%
Qwen2.5-VL-7B 2.2% 0.0% 0.0% 1.0% 3.0% 0.0% 0.0% 1.0%
OpenWebVoyagerIL 8.2% 5.9% 4.3% 6.3% 3.0% 5.8% 4.7% 6.6%
Web-CogReasoner (Ours) 16.3% 23.5% 15.2% 17.0% 12.1% 7.7% 9.3% 10.1%
The Mind2Web dataset (Table 7) serves as a crucial benchmark for evaluating generalization to unseen tasks (cross-task) and unseen websites (cross-web). Our model, Web-CogReasoner, records overall scores of 17.0% and 10.1% in these respective settings.
[…]
5.4 Ablation Study
Table 8: Ablation study of our Web-CogReasoner on the Web-CogBench. We evaluate the impact of each knowledge stage by progressively training the base model. "Base Model" is the original Qwen2.5-VL-7B. Each subsequent stage adds a new layer of knowledge training.
Model Configuration Memorizing Understanding Exploring Overall
Qwen2.5-VL-7B (Base Model) 67.6 61.0 77.9 69.8
+ Factual Knowledge (S1) 85.5 (+16.9) 64.2 60.1 72.1
+ Conceptual Knowledge (S2) 88.1 75.5 (+11.3) 65.8 78.3
+ Procedural Knowledge (S3) 90.8 74.1 85.0 (+7.1) 84.4
To empirically validate the effectiveness of our hierarchical, Bloom’s Taxonomy-inspired training methodology, we conduct a thorough ablation study. Our goal is to isolate and quantify the contribution of each distinct knowledge layer: Factual, Conceptual, and Procedural. We progressively build our Web-CogReasoner agent, starting from the base Qwen2.5-VL-7B model and sequentially adding each knowledge stage. The performance at each step is evaluated on our Web-CogBench, as its cognitive dimensions (Memorizing, Understanding, Exploring) directly align with our knowledge framework. The results are presented in Table 8.
Impact of Factual Knowledge
The first stage of our curriculum introduces Factual Knowledge, teaching the agent the "what" of web elements. As shown in Table 8, this initial stage provides a substantial performance boost, raising the overall score from 66.9% to 72.1%. Critically, the improvement is most pronounced in the Memorizing dimension, which surges from 68.6% to 85.5%. This demonstrates that by learning to perceive and identify concrete, factual information (e.g., element roles and names), the agent builds a solid perceptual foundation, which is the first and essential step towards more complex reasoning.
Impact of Conceptual Knowledge
Building upon this factual base, the second stage integrates Conceptual Knowledge, which focuses on the "how things relate." This layer enables the agent to understand the relationships between elements and the underlying semantic structure of a webpage. The results show another significant leap in performance, with the overall score climbing to 78.3%. The most notable gain occurs in the Understanding dimension, which improves by 11.3%. This confirms our hypothesis that once an agent can perceive facts, it must then learn to comprehend their context and purpose to effectively interpret a webpage’s functionality.
Impact of Procedural Knowledge
The final stage incorporates Procedural Knowledge, teaching the agent "how to act" to achieve a goal. This layer completes the Web-CogReasoner’s cognitive architecture, enabling it to formulate and execute plans. This stage delivers the final and most critical improvements for task completion, pushing the overall score to our final state-of-the-art result of 85.0%. The most dramatic increase is observed in the Exploring dimension, which skyrockets by 7.1%. This finding strongly indicates that goal-oriented, procedural training is paramount for converting passive understanding into active, successful task execution.
In summary, this staged analysis provides compelling empirical evidence for our core thesis. Each knowledge layer, from Factual to Conceptual to Procedural, is an indispensable component that progressively builds the agent’s cognitive capabilities, validating our structured, curriculum-based approach to training highly capable web agents.
5.5 Qualitative Analysis
Beyond the quantitative improvements shown in our ablation study, a qualitative analysis of the agent’s behavior at each stage offers deeper insights into how our curriculum shapes its cognitive abilities. We examine the agents’ performance on a representative task: "Find and add a laptop under \$1000 to the shopping cart on an e-commerce website."
Base Model (Qwen2.5-VL-7B)
Without any specialized training, the base model struggles to formulate a coherent plan. Its reasoning is often generic and untethered from the specific UI. It might correctly identify a "search bar" but fails to execute a meaningful action, or hallucinates actions that are not possible. For instance, its thought process might be: "I should search for a laptop," but its action is an ungrounded ‘click "Categories"‘ because it lacks the procedural knowledge to connect intent to a multi-step sequence of actions.
Stage 1 Agent (+ Factual Knowledge)
After training on Factual Knowledge, the agent’s perceptual abilities are significantly enhanced. It can now accurately identify and label key elements with their correct attributes. Its thought process becomes grounded in the facts of the page: "I see a search bar [ID: 25] with the name ’Search products’. I see a button [ID: 28] with the name ’Search’." However, it still struggles with planning. It understands "what" is on the page but not "why" or "how" to use it. It might correctly type "laptop" into the search bar but then get stuck, not understanding that the next logical step is to click the search button to submit the query.
Stage 2 Agent (+ Conceptual Knowledge)
With the addition of Conceptual Knowledge, the agent begins to understand the relationships between elements and their purpose. Its reasoning graduates from simple identification to semantic interpretation. The thought process now reflects this understanding: "The search bar [ID: 25] is for inputting queries. The search button [ID: 28] is functionally linked to it and will trigger the search. This group of elements forms a ’search component’." This allows it to reliably complete the search and navigate to the results page. However, on the results page, it may still struggle with complex procedural logic, such as applying a price filter.
Full Model (Web-CogReasoner, + Procedural Knowledge)
The final agent, equipped with Procedural Knowledge, demonstrates goal-oriented planning and execution. It seamlessly translates the high-level task into a concrete action sequence. Its thought process is now a strategic plan: "Goal: Add laptop under \$1000 to cart. Step 1: Type ’laptop’ into search bar [ID: 25]. Step 2: Click search button [ID: 28]. Step 3: On the results page, locate the ’Price Range’ filter. Step 4: Input ’1000’ into the ’max price’ field [ID: 57]. Step 5: Identify a suitable product from the filtered list and click its ’Add to Cart’ button [ID: 83]." This demonstrates a complete cognitive loop from perception and understanding to successful action, validating the necessity of the final procedural training stage.
Appendix A Appendix
A.1 Bloom’s Taxonomy and Web Cognitive
A.1.1 Bloom’s Tax
  • 1.
    Factual Knowledge: The foundational layer, encompassing the basic, discrete elements of a discipline that a student must know, such as essential terminology and specific, isolated details.
  • 2.
    Conceptual Knowledge: The synthesis of factual elements into a coherent, organized structure. This level focuses on the interrelationships between basic elements, including knowledge of classifications, principles, generalizations, theories, and models.
  • 3.
    Procedural Knowledge: The knowledge of how to perform a task or inquiry. This involves an understanding of specific skills, algorithms, techniques, and methods, representing a shift from "knowing-what" to "knowing-how."
A.1.2 Web data introduction
Cognitive science typically classifies human knowledge into three categories—factual, conceptual, and procedural—which correspond to different stages of cognitive development: perceiving, understanding, and executing. This taxonomy captures the natural trajectory of human learning: starting with the perception of concrete facts and data (factual knowledge), progressing toward abstract comprehension of concepts and relationships (conceptual knowledge), and ultimately acquiring the ability to carry out complex, goal-oriented behaviors through practiced routines and strategies (procedural knowledge). An illustrative example is shown in Figure5).
Figure 5: How people handle unfamiliar web pages. People learn factual, conceptual, and procedural knowledge to memorize, understand, and explore the web, ultimately completing specific tasks.
A.1.3 Data Sourcing
To collect comprehensive metadata from web pages, we developed a data collection tool based on Playwright. This tool performs deep traversal and interaction by systematically clicking on all elements within each page. We define each round of interaction (i.e., one click) as a layer, and using this iterative approach, we collected Layer 1 to Layer 6 data from 14 different websites (for complete website information, refer to Table 2 )
Table 9: Web elements’s meta-data.
Data Description
css element‘s CSS selectors
allcss CSS selector sequence of preceding elements
ourterhtml element’s outerhtml
location element’s boundingbox
role element’s role
name element’s name
Data precessing
For each web element, we capture its standalone screenshot, as well as screenshots taken before clicking (both with and without a red bounding box), after hovering, and after clicking. See Figure 6 for an example. We also collect the following metadata: CSS, allCSS, outerHTML, and location. Additionally, we extract semantic information from each element based on its outerHTML. If a role attribute is explicitly defined, we use its value directly as the element’s semantic role. Otherwise, we infer the role by mapping the tag name using the WAI-ARIA specification. Similarly, to determine the element’s semantic name, we extract the value of the aria-label if present; otherwise, we get its textual content. See Table 9 for detailed metadata of the web elements.
Figure 6: Example visual states of a web element (“USD”) we captured. Shown are: the element highlighted in the full-page view (base_rect.png), the hover state (hover.png), and the click state (click.png). These screenshots illustrate how the element’s visual context evolves through user interactions.
After collecting both the visual and semantic metadata, we present the corresponding screenshots of each clickable element to Qwen-VL 72B. The model is instructed to:
  • analyze the visual changes on the page after hovering over and clicking the target element;
  • identify and list any sub-elements that appear upon interaction (e.g., when a dropdown menu is triggered by clicking);
  • infer and generalize the functional purpose of the element.
For the functional purpose prediction, the model is additionally required to provide a confidence score. If this score remains below 0.5 after three retries, the prediction is excluded from evaluation.
A.2 Task definition
A.2.1 Factual Web Knowledge
Element Attribute Rec
We define the Element Attribute Recognition task to assess a model’s capability to infer the interactive semantics of web elements exclusively from visual input. Given a full-page screenshot with a specific interactive element marked by a red bounding box, the model is tasked with predicting two key attributes:
  • the semantic role of the highlighted element (e.g., "button", "link", "checkbox"),
  • the semantic name, which refers to the element’s accessible textual description (e.g., "Submit", "Search", "Next").
The ground truth for both attributes is derived from the element-level metadata collected as described in Section A.1.3. This task simulates the human cognitive ability to interpret the function of web interface components through visual perception alone, without relying on HTML structure or programmatic representations.
Sub-elements Pre
We define the Element Sub-element Prediction task to evaluate a model’s ability to infer the hierarchical structure of web interfaces—specifically, to identify the sub-elements that become visible upon interaction with a given parent element, using only visual information. In each task instance, the model is presented with a full-page screenshot in which a specific interactive element is highlighted by a red bounding box. The model is instructed to predict the set of sub-elements (e.g., menu items, dropdown options) that appear as a direct result of interacting with the highlighted element, such as clicking or hovering.The ground truth annotations for sub-elements are derived from the element-level metadata collected during the dynamic interaction process, as detailed in SectionA.1.3. This task simulates the human cognitive process of understanding interactive dependencies in a graphical interface—recognizing not only that a component is clickable, but also predicting its dynamic expansion behavior.
Page Change Pre
We define the Page Change Prediction task to evaluate a model’s capability to infer the visual consequences of interacting with a specific web element, relying solely on visual input. In this task, the model is presented with a full-page screenshot in which a target interactive element is highlighted by a red bounding box. The model is required to in an open-ended format to predict the visual changes that are likely to occur on the page after the element is clicked.The ground truth for this task is obtained from the generated responses of Qwen-VL 72B, which were produced based on visual metadata, as detailed in SectionA.1.3. This task is designed to simulate the human cognitive ability to anticipate the dynamic behavior of web interfaces through perception alone—without access to the underlying source code or prior knowledge of the page logic.
Next Page Pre
We define the Next Page Prediction task to evaluate a model’s ability to forecast navigation outcomes. Given a full-page screenshot with a highlighted interactive element, the model must predict the subsequent page that would result from interacting with that element. To ensure generalization capability, we implement two evaluation formats: multiple-choice selection (choosing from 4-5 possible next pages) and open-ended generation (describing the expected next page). Ground truth is derived from actual navigation sequences recorded during web interactions. This task assesses the agent’s understanding of functional relationships between interface elements and destination pages.
Source element Pre
We define the Source Element Prediction task to assess a model’s ability to identify which element on a webpage leads to a specific target page, using only visual input. The model is given two screenshots: one showing the current webpage with 4-10 candidate elements marked by bounding boxes, and another showing the resulting target page. Based on visual cues alone, the model should determine which candidate element would trigger the transition to the target page when interacted with. This task simulates the human ability to reason about visual cause-and-effect relationships in web navigation, without relying on code or prior knowledge of page logic.
A.2.2 Understanding Web Knowledge
Element Understanding
This task requires the model to produce a comprehensive, open-ended description of a highlighted element’s visual appearance, functional semantics, and placement on the webpage. Specifically, the output should cover: (1) Visual Traits (text, shape, iconography); (2) Location (e.g., top-right, footer); and (3) Function (e.g., navigates to user profile). This task simulates abstract comprehension from concrete element appearance.
WebPage Understanding
In this task, the model must generate a detailed and structured overview of the entire page. The response includes layout segmentation (e.g., header, sidebar, content area), key modules (e.g., search panel, product gallery), and a summary of page purpose and interactivity. This facilitates understanding of page-wide structure and intent.
Caption & QA
We define the Caption & QA task to evaluate a model’s capability to comprehend and reason over both image and non-image content embedded within webpages. This task comprises four subtasks:
  • Embedded Image Captioning: Given a full-page screenshot containing one or more embedded images, the model is required to generate a detailed and semantically meaningful caption for each image, describing its visual content and its contextual relevance within the surrounding webpage layout.
  • Embedded Image QA: Given a question grounded in the content of an embedded image within a webpage screenshot, the model must produce an accurate, context-aware answer using only visual information. These questions may refer to image content (e.g., "What brand is shown in the ad?") or its function in the UI.
  • Webpage Captioning: The model is tasked with generating an open-ended description of the webpage’s content, layout, and interactive purpose, treating the entire screenshot as input. The generated caption should reflect both structural composition and the inferred user intent of the webpage.
  • Webpage QA: Given a full-page screenshot and a natural language question referring to any aspect of the page (e.g., title, layout, purpose, textual content), the model must generate a grounded and precise answer based on visual and spatial information.
All four subtasks are derived from the Multi-UI [23] dataset, which provides rich annotations for webpage visual elements and user-facing semantics. Together, these subtasks measure a model’s ability to perform grounded visual-language understanding at both local (element-level) and global (page-level) scales.
A.2.3 Procedual Web Knowledge
User’s intention Pre
Built on the MultiModal-Mind2Web dataset—which provides natural language instructions, action trajectories, and aligned web page screenshots—we introduce a novel multi-modal task: inferring the user’s high-level intent from a sequence of visual observations. Unlike traditional imitation learning or instruction-following tasks, our setting requires the model to infer why a trajectory occurred, rather than how to execute it. Solving this task demands both visual understanding and temporal reasoning. The details of this task are as follows:
  1. 1.
    Task Definition: Given a sequence of web page screenshots $p_{1},p_{2},\ldots,p_{n}$ representing a user’s interaction trajectory, the objective is to predict the original user instruction y that guided the sequence. Each screenshot $p_{t}$ corresponds to the visual observation at step $t$ of a successful task execution. Formally, the model learns a mapping: $f:\{p_{1},p_{2},……,p_{n}\}\rightarrow Q$ where Q is the natural language instruction.
  1. 2.
    Dataset Construction: We construct our dataset by processing the original MultiModal-Mind2Web corpus. For each task, we extract only the visual observations—i.e., the sequence of web page screenshots corresponding to each step in the execution trajectory. We then pair each screenshot sequence with the original natural language instruction as the supervision signal.
Popup Close
We curated a collection of 51 popup components from JS Design website, encompassing a diverse range of visual styles and functional categories, such as notification modals, alert dialogs, and login forms. This diversity ensures comprehensive coverage of real-world popup use cases. For background webpages, we utilized the OpenWebVoyager[13] dataset, which contains a large number of authentic webpage screenshots with varied layouts and content, providing a rich foundation for synthesizing realistic popup-injected webpages. To construct the training data for this task, we employed the following procedure:
  1. 1.
    Synthesizing Webpage Screenshots with Popups: We randomly overlaid popup images onto background webpage screenshots to simulate webpages containing popups. During synthesis, we introduced variability by randomly adjusting the popup’s size and position and modifying the brightness and sharpness of the background images, thereby enhancing visual diversity and realism.
  1. 2.
    Generating Popup AX Tree: Each popup image was processed using Qwen-VL 2.5 32B to generate an ARIA-compliant AX Tree. To simulate diverse structural configurations, we randomly modified the index values of popup AX Tree elements and inserted the popup AX Tree into different locations within the original webpage’s AX Tree, resulting in a combined AX Tree that reflects realistic variations in webpage structure.
  1. 3.
    Generating Popup Closing Strategies: We then instructed Qwen-VL 2.5 32B to identify all n possible methods for closing the popup, based on the popup image and its corresponding AX Tree. Recognizing that, in practical settings, any correct method is sufficient, we applied combinatorial augmentation to the n methods. Specifically, we enumerated all non-empty subsets of the n strategies, yielding a total of $2^{n}-1$ distinct answer combinations. This expansion significantly broadens the training distribution and increases the model’s exposure to diverse correct solutions.
  1. 4.
    Constructing the Training Dataset: Using the synthesized webpage screenshots and the enriched AX Tree, we constructed a dataset for training models on popup dismissal. Each data point comprises:
    • Input: a webpage screenshot with an embedded popup and the corresponding AX Tree;
    • Output: valid methods for closing the popup.
Single-Step Web Task
We define the Single-Step Web Task to evaluate a model’s ability to ground high-level user intentions in visual webpage elements. Each task instance includes a full-page screenshot from a real-world webpage, a concise natural language instruction (e.g., "Search for a product", "Log into the system"), and several candidate elements marked by red bounding boxes.
The model must identify which element, if clicked, would successfully fulfill the given task. This setup simulates perceptual grounding of user intent—matching natural language goals to actionable UI targets based solely on visual cues.
All samples are directly sourced from the Multi-UI [23] dataset, which provides rich, annotated webpage screenshots paired with task descriptions and labeled ground-truth targets. No trajectory-level annotation or external instruction rewriting is involved. This task offers a reliable benchmark for evaluating atomic web interaction capabilities in a static, visually grounded setting.
Noisy Multi Step Web Task
To further enhance the original OpenWebVoyager [13] dataset, we incorporate Knowledge-driven Chain-of-Thought (CoT) Reasoning to improve the model’s stepwise understanding and execution, details see Section  4.2 In addition, to simulate realistic interruptions during multi-step web interactions, we propose the Noisy Multi-Step Web Task by augmenting interaction trajectories from OpenWebVoyager. Specifically, for each sample in our Popup Close dataset, a popup window is injected at a specific step (e.g., step $t$) of an existing task trajectory.
This modification introduces a prerequisite interaction: the agent must first detect and dismiss the popup before resuming progress toward the original task goal. By explicitly modeling such interruptive UI elements, this task formulation captures a more realistic web interaction paradigm in which user flows are frequently obstructed. It also provides a challenging benchmark for evaluating agents’ robustness to UI-level noise and their capacity for error recovery.

### v2

@@
 In this paper, we argue that web agents must first acquire sufficient knowledge to effectively engage in cognitive reasoning [^1] Drawing inspiration from Bloom’s educational philosophy, a cornerstone of modern pedagogy..
-Therefore, we decompose a web agent’s capabilities into two essential stages: knowledge content learning and cognitive processes.
-To formalize this, we propose Web-CogKnowledge Framework, which categorizes knowledge into Factual, Conceptual, and Procedural domains.
-In this framework, knowledge content learning corresponds to the agent’s processes of Memorizing and Understanding, which rely on the former two types knowledge respectively, representing the "what" of learning.
-Conversely, cognitive processes correspond to Exploring, grounded in Procedural knowledge, defining the "how" of reasoning and action.
-To facilitate knowledge acquisition, we construct the Web-CogDataset, a structured resource curated from 14 real-world websites, designed to systematically instill the core knowledge necessary for web agent.
-This dataset serves as the agent’s conceptual grounding—the "nouns" upon which comprehension is built—as well as the basis for learning how to reason and act.
-Building on this foundation, we operationalize these processes through a novel knowledge-driven Chain-of-Thought (CoT) reasoning framework, developing and training our proposed agent, the Web-CogReasoner.
+Therefore, we decompose a web agent’s capabilities into two essential phases: knowledge content learning and cognitive processes.
+To formalize this, we propose the Web-CogKnowledge Framework, which categorizes knowledge into Factual, Conceptual, and Procedural domains.
+In this framework, knowledge content learning corresponds to the agent’s processes of Memorizing and Understanding, which rely on the first two types of knowledge respectively, representing the ”what” of learning.
+Conversely, cognitive processes correspond to Exploring, grounded in Procedural knowledge, defining the ”how” of reasoning and action.
+To facilitate knowledge acquisition, we construct Web-CogDataset, a structured resource curated from 14 real-world websites, designed to systematically instill the core knowledge necessary for a web agent.
+This dataset serves as the agent’s conceptual grounding—the ”nouns” upon which comprehension is built—as well as the basis for learning how to reason and act.
+Building on this foundation, we operationalize these processes through a novel knowledge-driven Chain-of-Thought (CoT) reasoning framework, developing and training our proposed agent, Web-CogReasoner.
 Extensive experimentation reveals its significant superiority over existing models, particularly in its capacity for generalization to unseen tasks where its structured knowledge proves decisive.
-To facilitate rigorous and systematic evaluation, we introduce the Web-CogBench, a comprehensive evaluation suite designed to assess and compare agent performance across the delineated knowledge domains and cognitive capabilities.
+To facilitate rigorous and systematic evaluation, we introduce Web-CogBench, a comprehensive evaluation suite designed to assess and compare agent performance across the delineated knowledge domains and cognitive capabilities.
 Our code and data is open sourced at https://github.com/Gnonymous/Web-CogReasoner.
@@
 1 Introduction
-The advent of large-scale models marks a landmark advancement in artificial intelligence.
-The development of Large Multimodal Models (LMMs), in particular, has vastly expanded the horizons for AI applications.
-In recent years, AI agents have emerged as a primary vehicle for deploying these models, significantly extending their capabilities.
-Much like humans, these agents can leverage a diverse array of tools to enhance their skills, leading to remarkable achievements in domains such as code generation hui2024qwen2 ; jiang2024survey , image huang2025diffusion ; bie2024renaissance and video synthesis  assran2025v ; liu2024sora , and academic research.
-Notably, the application of web agents to surfing the internet has also seen substantial progress.
-The evolution of web agents has transitioned from early rule-based systems to modern approaches that widely leverage Large Language Models (LLMs) or Languge Vision Models (LVMs) wang2024survey ; ning2025survey ; zhang2024large ; sapkota2025ai .
-These LLM-powered agents typically process HTML or Accessibility Tree inputs, formulating tasks and observations into natural language prompts for the model to reason upon and decide the next action.
-With the integration of LVMs, web agents have gained perceptual abilities akin to human vision, enabling them to comprehend and reason about multimodal content on web pages.
-As a whole, web agents can be broadly categorized into three types: (1) Text-only agents zhou2023webarena ; li2023camel , which suffer from information loss due to their inability to capture visual cues; (2) Vision-only agents qin2025ui , which rely solely on visual information and thus lose critical structured data; and (3) Hybrid agents koh2024visualwebarena ; he2024openwebvoyager , which utilize both visual and textual information.
-LLMs and LMMs are pre-trained on vast amounts of general-domain knowledge, which is undeniably crucial for their foundational capabilities.
-However, as with any specialized field, this general knowledge is often insufficient, creating a performance bottleneck that caps the potential of the three aforementioned model types.
-While many existing works have attempted knowledge enhancement or injection, these methods often lack a systematic or theoretically grounded approach.
-To address this, we draw inspiration from a highly successful paradigm in human education: Bloom’s Taxonomy ormell1974bloom ; conklin2005taxonomy .
-This framework suggests that effective learning can be deconstructed into two key stages: knowledge content learning and cognitive processes.
-Accordingly, we propose that training a web agent should follow a similar two-stage methodology.
-The first stage is dedicated to acquiring knowledge of varying difficulties.
-The second stage then focuses on the analysis, application, and ultimately, the creative synthesis of that knowledge.
-This mirrors the human learning trajectory: we first accumulate knowledge through education (Stage 1), and then based on that foundation of knowledge and experience, we learn to apply, innovate, and create (Stage 2).
-Therefore, in this paper, we introduce a training paradigm for our model grounded in Bloom’s Taxonomy, bifurcated into two sequential stages: Knowledge Content Learning and Cognitive Processes.
-The initial stage, Knowledge Content Learning, is dedicated to building a multi-layered knowledge foundation.
-This foundation consists of three distinct types: Factual knowledge, which includes basic, foundational concepts; Conceptual knowledge, which maps the interrelationships between these facts; and Procedural knowledge, which represents the logical frameworks needed to synthesize factual and conceptual understanding for effective reasoning.
-To curate this knowledge, we construct Web-CogDataset, collecting metadata from 14 prominent websites and engineered 12 granular tasks designed to systematically instill each knowledge type.
-In essence, this stage teaches the agent foundational web knowledge and how to apply it to familiar tasks.
-The subsequent stage, Cognitive Processes, focuses on cultivating the agent’s self-exploring capability.
-This advanced capability empowers the agent to build upon its learned foundation, enabling it to "learn how to learn" by creatively exploring its own "self-knowledge" to tackle previously unencountered tasks and challenges.
-To operationalize this, we developed knowledge-guided reasoning templates and utilized imitation learning to train the model.
-This process explicitly teaches the agent the cognitive faculties required for autonomous learning and creative problem-solving in novel scenarios.
-To validate our approach, we conducted rigorous evaluations on a suite of general-purpose benchmarks as well as our own.
-Our findings reveal that the agent trained with our methodology consistently and significantly outperforms state-of-the-art baselines.
-Crucially, and in alignment with our hypothesis, the performance gap widens dramatically on knowledge-intensive tasks.
-This demonstrates that, analogous to human learning, the structured acquisition of knowledge is a key determinant of an agent’s ability to excel in complex, domain-specific scenarios.
+The advent of large-scale models marks a milestone in artificial intelligence, with Large Multimodal Models (LMMs) greatly expanding application horizons.
+AI agents have become the primary vehicle for deploying these models, enabling capabilities in code generation (Hui et al., 2024; Jiang et al., 2024), image and video synthesis (Huang et al., 2025; Bie et al., 2024; Assran et al., 2025; Liu et al., 2024c), and academic research (Li et al., 2025; Zhang et al., 2025).
+Recent progress has also highlighted the growing importance of web agents.
+Web agents have evolved from early rule-based systems to modern approaches leveraging Large Language Models (LLMs) and Language Vision Models (LVMs) (Wang et al., 2024; Ning et al., 2025; Zhang et al., 2024; Sapkota et al., 2025).
+LLM-powered agents typically convert HTML or Accessibility Tree inputs into natural language prompts for reasoning and action.
+With LVMs, agents gain perceptual abilities akin to human vision, allowing them to process multimodal content on web pages.
+Broadly, web agents can be categorized as: (1) text-only (Zhou et al., 2023; Li et al., 2023), which miss visual cues; (2) vision-only (Qin et al., 2025), which lack structured data; and (3) hybrid (Koh et al., 2024; He et al., 2024b), which integrate both modalities.
+LLMs and LVMs pre-trained on general-domain knowledge provide strong foundations but remain limited in specialized tasks, creating performance bottlenecks.
+Prior knowledge-enhancement methods often lack systematic or theoretical grounding.
+To address this, we draw inspiration from Bloom’s Taxonomy (Ormell, 1974; Conklin, 2005), which divides learning into two phases: Knowledge Content Learning and Cognitive Processes.
+In our paradigm, the first phase builds a multi-layered foundation: Factual Knowledge, covering basic concepts, and Conceptual Knowledge, capturing their interrelations.
+This equips the agent with core web knowledge and its application to familiar tasks.
+The second phase develops Procedural Knowledge, providing logical frameworks to synthesize prior knowledge for reasoning and exploration.
+This enables the agent to “learn how to learn,” creatively leveraging self-knowledge to solve novel challenges.
+This mirrors the human learning trajectory: we first accumulate knowledge through education (Phase 1), and then based on that foundation of knowledge and experience, we learn to apply, innovate, and create (Phase 2).
+To support this, we construct Web-CogDataset from 14 prominent websites with 12 fine-grained tasks, and design knowledge-guided reasoning templates combined with imitation learning to instill the required cognitive faculties.
+Rigorous evaluations on public and in-domain benchmarks show that our method consistently surpasses state-of-the-art baselines, with the performance advantage especially pronounced in knowledge-intensive tasks.
+These results confirm that structured knowledge acquisition is crucial for enabling agents to excel in complex, domain-specific scenarios.
 In summary, our contributions are threefold:
   1. 1.
-    Drawing inspiration from Bloom’s taxonomy and established human educational paradigms, we propose the Web-CogKnowledge Framework — a systematic, two-stage training methodology designed to enhance the cognitive capabilities of web agents.
-Built upon this framework, we develop Web-CogReasoner.
+    Drawing inspiration from Bloom’s taxonomy and established human educational paradigms, we propose the Web-CogKnowledge Framework, a systematic, two-phase training methodology designed to enhance the cognitive capabilities of web agents.
+As shown in Figure 1, built upon this framework, we develop Web-CogReasoner.
 Rigorous benchmarking demonstrates that agents trained under our framework achieve a significant performance improvement over current state-of-the-art models.
   2. 2.
-    We construct the Web-CogDataset, a structured curriculum consisting of 12 fine-grained and progressively challenging tasks.
+    We construct Web-CogDataset, a structured curriculum consisting of 12 fine-grained and progressively challenging tasks.
 These tasks are meticulously designed to incrementally build the agent’s web knowledge, cognition capability, and higher-order reasoning.
@@
 2 Related Work
-Web Agent
-Early studies on web page understanding have concentrated on leveraging structured text.
-The work gur2022understanding defined crucial HTML understanding tasks, such as semantic classification, description generation, and autonomous web navigation, and trained models to comprehend the structure and content of HTML.
-AutoWebGLMlai2024autowebglm introduced an HTML simplification algorithm.
-Through curriculum learning, it first trained the model using collected web page data to recognize web page structures and understand the functions of different interactive components, like text boxes and buttons.
-Subsequently, the model learned to execute simple web tasks and decompose complex ones into subtasks, enabling the complete execution of web tasks from simple to complex vision-based Web Page Understanding.
-Recent research has explored the use of visual information for web page understanding.
-SeeClickcheng2024seeclick designed specialized pre-training tasks, including text-to-point and text-to-bounding box localization tasks, as well as point-to-text and bbox-to-text generation tasks.
-These tasks facilitated the model’s learning of the correlation between the positions of web elements and their textual descriptions, enhancing the model’s ability to localize and understand web elements.
-CogAgenthong2024cogagent introduced a high-resolution cross-module, enabling the model to recognize tiny elements and text in GUIs, and constructed a large-scale GUI annotated dataset for GUI grounding tasks, achieving excellent performance in multiple visual question answering (VQA) benchmarks and GUI navigation task datasets.
-OmniParserwan2024omniparser integrated three typical visually-situated text parsing tasks—text spotting, key information extraction, and table recognition—within a unified framework, providing a more efficient solution for GUI Agents to handle complex tasks.
-Unlike many existing agent frameworks that rely on heavily wrapped commercial models (e.g., GPT-4o) with expert-crafted prompts and workflows, UI-TARSqin2025ui adopted an end-to-end architectural design, directly perceiving screenshots, executing reasoning processes, and autonomously generating valid actions.
-UGroundgou2024navigating developed a universal visual grounding model using a GUI dataset with 10 million elements, demonstrating outstanding performance in desktop tasks and mobile tasks.
-Multimodal approaches have emerged as a powerful way to comprehensively understand web pages.
-WebVoyagerhe2024webvoyager utilized the GPT-4V-ACT tool to overlay bounding boxes and numerical labels of interactive elements on web page screenshots.
-Additionally, it complemented web page visual information with the Accessibility Tree, helping the model gain a more comprehensive understanding of web page content.
-SeeActzheng2024gpt directly harnessed the powerful visual and language understanding capabilities of GPT-4V and explored various grounding strategies to map the text plans generated by the model into specific web operations.
-Tuning-free Attention-driven Grounding (TAG) xu2025attention method further accomplishes this task without the need for additional fine-tuning by leveraging the inherent attention patterns in pretrained MLLMs.
-Web Agent Evaluation
-To evaluate GUI agents’capabilities in web interactions, benchmarks are broadly categorized into web browsing tasks and web understanding assessments.
-For web browsing, offline benchmarks like Mind2Webdeng2023mind2web , Multimodal-Mind2Web, AutoWebBenchlai2024autowebglm and WebVLN-v1chen2024webvln evaluate models’ability to perform multi-step web tasks (e.g., booking flights, filling dynamic forms) using static datasets that include webpage content along with corresponding interaction trajectories.
-In contrast to offline benchmarks, online environments such as Mini-WoB++liu2018reinforcement , Webshopyao2022webshop , and WebArenazhou2023webarena enable real-time interaction testing.
-Mini-WoB++liu2018reinforcement focuses on low-level UI operations (e.g., clicking buttons, handling pop-ups) within simplified webpages.
-By comparison, Webshopyao2022webshop and WebArenazhou2023webarena replicate complex, real-world tasks—such as e-commerce transactions and Wikipedia editing—to evaluate task completion performance.
-Building on this, VisualWebArenakoh2024visualwebarena incorporates multimodal inputs to assess models’ ability to handle vision-augmented interactions in dynamic web environments.
-To evaluate deeper web understanding, benchmarks are designed to target specific cognitive abilities.
-WEBQAchang2022webqa assesses a model’s capacity for open-domain, multi-hop reasoning.
-In contrast, ScreenQAhsiao2022screenqa and the ScreenAIbaechler2024screenai suite focus on screen-centric comprehension.
-ScreenQAhsiao2022screenqa evaluates a model’s ability to recognize UI elements and interpret context by answering questions based on screenshots.
-ScreenAIhsiao2022screenqa further extends this with three progressively challenging subtasks: Screen Annotation (SA), which involves spatial layout tagging; ScreenQA Short, which tests concise fact extraction; and Complex ScreenQA, which requires multi-step reasoning across diverse screen layouts.
-Together, these tasks assess fine-grained layout understanding, semantic interpretation, and practical problem-solving in visually dense interfaces.
+2.1 Web Agent
+Early work on web understanding focused on structured HTML, addressing tasks like semantic classification, description generation, and navigation (Gur et al., 2022), with AutoWebGLM (Lai et al., 2024) further applying curriculum learning for structure recognition, component understanding, and progressively complex task execution.
+More recent studies leverage visual signals: SeeClick (Cheng et al., 2024) linked elements to textual descriptions to enhance localization, CogAgent (Hong et al., 2024) combined high-resolution cross-module modeling with a large GUI dataset for VQA and navigation, OmniParser (Wan et al., 2024) unified text spotting, extraction, and table recognition, UI-TARS (Qin et al., 2025) directly maps screenshots to actions, and UGround (Gou et al., 2024) trained on 10M GUI elements for robust desktop and mobile performance.
+Multimodal approaches integrate text and vision: WebVoyager (He et al., 2024a) combines screenshots with bounding boxes and accessibility trees, SeeAct (Zheng et al., 2024) grounds text plans via GPT-4V, and TAG (Xu et al., 2025) exploits pretrained attention for grounding without fine-tuning.
+2.2 Web Agent Evaluation
+Benchmarks are categorized into browsing and understanding.
+For browsing, offline datasets like Mind2Web (Deng et al., 2023), Multimodal-Mind2Web, AutoWebBench (Lai et al., 2024), and WebVLN-v1 (Chen et al., 2024) test multi-step task execution, while online environments such as Mini-WoB++ (Liu et al., 2018), Webshop (Yao et al., 2022), and WebArena (Zhou et al., 2023) allow real-time evaluation.
+Mini-WoB++ emphasizes low-level UI operations, whereas Webshop and WebArena simulate complex tasks.
+VisualWebArena (Koh et al., 2024) further adds multimodal inputs for dynamic interactions.
+For web understanding, WEBQA (Chang et al., 2022) evaluates open-domain multi-hop reasoning.
+ScreenQA (Hsiao et al., 2022) and ScreenAI (Baechler et al., 2024) focus on screen comprehension, with ScreenQA targeting UI recognition and contextual QA, while ScreenAI extends this into three subtasks: Screen Annotation, ScreenQA Short, and Complex ScreenQA.
+Together, they assess layout understanding, semantic interpretation, and reasoning in visually dense interfaces.
 3 Web-CogKnowledge Framework
 3.1 Bloom’s Taxonomy
-The Bloom’s Taxonomy  [^2] https://fctl.ucf.edu/teaching-resources/course-design/blooms-taxonomy/.
-(Anderson & Krathwohl, 2001) presents a two-dimensional pedagogical design, i.e., knowledge content learning and cognitive process dimension, which demonstrates a "shallow-to-deep" instructional methodology.
-This approach posits that effective learning is not monolithic but is a structured progression through distinct types of knowledge, beginning with foundational elements and culminating in high-level self-awareness.
-This educational strategy ensures that learners first establish a solid base of facts and concepts before advancing to the execution of complex procedures and, ultimately, the application of reflective metacognitive strategies.
-This pedagogical progression from simple to complex can be formally outlined through four hierarchical types of knowledge: Factual Knowledge, Conceptual Knowledge, Procedural Knowledge and Metacognitive Knowledge.
-More details can be found in Appendix A.1.1.
+The Bloom’s Taxonomy [^2] https://fctl.ucf.edu/teaching-resources/course-design/blooms-taxonomy/ (Anderson & Krathwohl, 2001) presents a two-dimensional framework: knowledge content learning and cognitive processes, embodying a ”shallow-to-deep” instructional methodology.
+It structures learning from foundational facts and concepts to complex procedures, ensuring a solid knowledge base before higher-level reasoning.
+This progression can be formalized through four hierarchical types of knowledge: Factual, Conceptual, and Procedural Knowledge. Further details are in Appendix A.1.1.
+Figure 2: Illustration of Web-CogDataset.
 3.2 Web-CogKnowledge
@@
   1. 1.
-    Factual knowledge: concrete information extracted from web content.
-Such as identifying the attributes of individual web elements and predicting the immediate, direct consequences of a single interaction;
+    Factual knowledge: concrete information extracted from web contents, such as identifying the attributes of individual web elements and predicting the immediate, direct consequences of a single interaction.
   2. 2.
-    Conceptual knowledge: semantic relationships and abstract patterns underlying webpage content and structure.
-Such as inferring the function of interface components, comprehending the overall purpose and structure of a webpage, and interpreting its multimodal content;
+    Conceptual knowledge: semantic relationships and abstract patterns underlying webpage content and structures, such as inferring the function of interface components, comprehending the overall purpose and structure of a webpage, and interpreting its multimodal content.
   3. 3.
     Procedural knowledge: actionable know-how for accomplishing specific tasks through interaction, including planning, decision-making, and sequential execution.
-Such as executing goal-oriented action sequences, inferring user intent from observed behaviors, and handling unexpected interruptions to complete complex tasks;
+Such as executing goal-oriented action sequences, inferring user intent from observed behaviors, and handling unexpected interruptions to complete complex tasks.
 This taxonomy aligns each knowledge type with a corresponding cognitive competency required for web-based reasoning and interaction.
-Figure 2: Statistics of Selected Websites by Category.
-To fully leverage the potential of Web-CogKnowledge, leading to the ultimate realization of the Web-CogKnowledge Framework.
-We first collect multimodal metadata from 14 real-world websites (see Figure 2) and then design Web-CogDataset with diverse set of web tasks for Web-Agent training in Section 3.3.
-And finally construct Web-CogBench in Section 3.4, built from meticulously chosen subsets.
-Specifically, we curate the metacognitive knowledge and Web-CogReasoner learning and training process in Section 4 due to its importance and complexity.
+To fully leverage the potential of Web-CogKnowledge, leading to the ultimate realization of the Web-CogKnowledge Framework, we first collect multimodal metadata from 14 real-world websites (see Figure 5) and then design Web-CogDataset with diverse sets of web tasks for web-agent training in Section 3.3, and finally construct Web-CogBench in Section 3.4, built from meticulously chosen subsets.
+Specifically, we curate Web-CogReasoner learning and training process in Section 4 due to its importance and complexity.
 More details about data collection and cleaning process can be found in Appendix A.1.5.
 3.3 Web-CogDataset
-By selectively crawling metadata from the aforementioned websites and leveraging the hierarchical structure of Web-CogKnowledge, we define a series of tasks corresponding to different levels of knowledge granularity.
-This process results in the construction of a comprehensive and hierarchically structured dataset, referred to as Web-CogDataset.
-The task definitions, organized according to the knowledge hierarchy, are introduced below.
-Detailed statistics and illustrative examples for Web-Dataset are provided in Table 1 and Figure 3.
-For additional implementation details, please refer to the Appendix A.2.
-Table 1: Statistics of the Web-CogDataset.
-Knowledge Task Statistics
-Subtotal Total
-Factual Web Knowledge Element Attribute Recognition 37K 81K
-Sub-elements Prediction 1K
-Page Change Prediction 24K
-Next Page Prediction 18K
-Source Element Prediction 1K
-Conceptual Web Knowledge Element Understanding 40K 62K
-WebPage Understanding 7K
-Caption & QA 15K
-Procedural Web Knowledge User’s Intention Prediction 2K 27K
-Popup Close 1K
-Single-Step Web Task 17K
-Noisy Multi-Steps Web Task 7K
-Factual Web Knowledge Tasks Overview
-The Factual Web Knowledge tasks are designed to establish the agent’s core perceptual abilities by operationalizing our definition of factual knowledge.
-These five tasks train the agent to identify the attributes of individual web elements and predict the immediate, direct consequences of a single interaction.
-By focusing on extracting concrete, visually grounded information—rather than abstract reasoning—this task category ensures the agent builds a foundational understanding of the web’s basic cause-and-effect mechanics.
-This layer of knowledge is the essential prerequisite for any subsequent conceptual or procedural reasoning.
-  • Element Attribute Recognition: Given a screenshot with a highlighted interactive element, the model predicts its semantic role (e.g., button, link) and accessible name (e.g., "Submit", "Search"), relying solely on visual cues.
-  • Sub-elements Prediction: The model infers the dynamic hierarchical structure of a web interface by predicting the set of sub-elements that appear upon interacting with a highlighted parent element.
-  • Page Change Prediction: The model anticipates the visual changes that occur after clicking on a specific element, formulating an open-ended response that describes the expected dynamic behavior of the page.
-  • Next Page Prediction: The model predicts the subsequent page that results from interacting with a specific element on the current page.
-To enhance generalization, we designed two types of tasks: multiple-choice questions and open-ended responses.
-  • Source Element Prediction: Given two screenshots—the current page and the resulting target page.
-Then the model identifies which of the visually marked elements on the current page leads to the target, simulating visual cause-and-effect reasoning.
-Figure 3: Illustration of Web-CogDataset.
-Conceptual Web Knowledge Tasks Overview
-Building upon the factual foundation, the Conceptual Web Knowledge tasks are designed to move the agent from perception to comprehension.
-These tasks train the agent to synthesize discrete facts into meaningful patterns and understand the semantic relationships underlying a webpage’s content and structure.
-Rather than merely identifying elements, the agent learns to infer their functions, comprehend the webpage’s overall purpose and organization, and interpret its multimodal content.
-This layer of conceptual understanding is critical, as it enables the agent to grasp the "why" behind the interface, transforming it from a collection of static elements into a coherent, interactive system.
-  • Element Understanding: For a specific interactive element, the model generates an open-ended paragraph that comprehensively describes the element’s Visible Traits (e.g., text, shape, styling), its On-page Location (e.g., header, sidebar, main content), and its likely User-facing Function (e.g., playing a video, navigating to a new page), relying solely on visual context.
-  • WebPage Understanding: Given a full-page screenshot, the model generates a comprehensive overview describing the webpage’s Layout Organization (e.g., header, event information, seating chart, filter panel), Key Element Analysis (e.g., element attribute, description, function, Interaction, expected outcome), and a Summary of the WebPage.
-This enables a thorough understanding of the webpage’s structure and functionality.
-  • Caption & QA: The model generates captions and answers questions about embedded images and WebPages, assessing its visual reasoning and holistic understanding of multimodal webpage information (e.g., layout, text).
-Contributes to further enhancing the model’s understanding of WebPages, while preserving its open conversational abilities with web content.
-The orignal data is drawn from the Multi-UI liu2024harnessing dataset.
-Procedural Web Knowledge Tasks Overview
-With both factual and conceptual knowledge established, the Procedural Web Knowledge tasks are designed to transition the agent from understanding to action.
-These tasks train the agent to formulate and execute goal-oriented plans by leveraging its actionable know-how for accomplishing specific tasks through interaction.
-The agent learns to decompose complex user goals into sequential steps, handle unexpected interruptions like popups, and orchestrate a series of interactions to achieve a desired outcome.
-This final knowledge layer is paramount, as it empowers the agent to apply its understanding strategically, converting passive knowledge into active, successful task completion.
-  • User’s Intention Prediction: The model infers high-level user intent from a sequence of webpage screenshots representing an interaction trajectory, requiring visual understanding and temporal reasoning.
-The task is built on the MultiModal-Mind2Web deng2023mind2web dataset, mapping screenshot sequences to natural language instructions.
-  • Popup Close: The model identifies and dismisses popups (e.g., notification modals, login forms) on synthesized webpage screenshots, using a dataset of 51 popup components from JS Design  [^3] https://js.design/.
-overlaid on OpenWebVoyager he2024openwebvoyager webpages, with combinatorial augmentation of closing strategies for diverse training.
-  • Single-Step Web Task: The model executes goal-directed web interactions to perform atomic web actions, using annotated screenshots from the MultiUI liu2024harnessing dataset, with tasks validated for accuracy.
-  • Noisy Multi-Steps Web Task: The model perform complex, goal-oriented interactions across multiple steps, based on unified OpenWebVoyager he2024openwebvoyager trajectories, with a standardized action space for consistent evaluation.
-To test robustness, some tasks include UI interruptions (e.g., popups) that must be dismissed before resuming the task.
-Table 2: Detailed statistics of Web-CogBench.
+By selectively crawling metadata from 14 representative websites and aligning it with the hierarchical design of Web-CogKnowledge, we construct Web-CogDataset, a large-scale, hierarchically structured dataset tailored for knowledge-centric web reasoning.
+The dataset spans three layers of knowledge: Factual, Conceptual, and Procedural.
+Each mapped to carefully designed task families that progressively cultivate perception, comprehension, and action-oriented reasoning.
+As illustrated in Figure 2, these tasks together form a coherent pipeline that transitions agents from identifying elemental attributes, to grasping semantic patterns and page structures, and finally to executing complex, goal-directed interactions under realistic constraints.
+This organization mirrors human learning trajectories, ensuring that higher-order reasoning is built on solid perceptual and conceptual foundations.
+Detailed task definitions, implementation protocols, and examples are provided in Appendix A.2.
+Dataset statistics are reported in Table 13.
+To ensure data quality and diversity, we further analyze the dataset composition and annotation reliability in Appendix A.1.4.
+3.4 Web-CogBench
+To evaluate the cognitive capabilities enabled by our knowledge-centric framework, we introduce Web-CogBench.
+While our training dataset is organized by knowledge type (Factual, Conceptual, Procedural), Web-CogBench measures agent performance across three corresponding abilities: Memorizing, Understanding, and Exploring.
+Curated from a representative subset of Web-CogDataset, it assesses how effectively an agent applies learned knowledge in complex web contexts.
+Detailed statistics are in Table 1, and the evaluation dimensions align with our hierarchical knowledge framework.
+A complete definition of all tasks is provided in the appendix A.2.2.
+Table 1: Statistics of Web-CogBench.
 Task Congition Metric #Num
@@
 Total - - 876
-3.4 Web-CogBench
-To rigorously evaluate the cognitive capabilities fostered by our knowledge-centric framework, we introduce Web-CogBench.
-While our training dataset is structured by knowledge types (Factual, Conceptual, Procedural), Web-CogBench assesses agent performance through the lens of three corresponding cognitive abilities: Memorizing, Understanding, and Exploring.
-This benchmark is curated from a representative subset of our Web-CogDataset and is designed to measure how well an agent translates its learned knowledge into practical skills within complex web contexts.
-Detailed statistics of Web-CogBench are provided in Table 2.
-The three evaluation dimensions directly mirror our hierarchical knowledge framework:
 Memorizing
@@
 4.1 Problem Setup
-We model the interaction between the environment and Web-CogReasoner as a partially observable Markov decision process (POMDP): $P=(S,A,O,K,T,R)$, where $S$ denotes the webpage status, $A$ is the action space (See complete operations in Table 1).
-$O$ is the observation space, and $K$ represents the internal knowledge of Web-CogReasoner.
-$T:S\times A\rightarrow S$ is the transition function.
+We model the interaction between Web-CogReasoner and the environment as a partially observable Markov decision process (POMDP): $P=(S,A,O,K,T,R)$, where $S$ denotes the webpage state, $A$ the action space (Table 10), $O$ the observation space, $K$ the internal knowledge, $T$ the transition function, and $R$ the reward function.
+At each step $t$, the agent receives a screenshot and its accessibility tree (AX Tree), forms a reasoning thought $h_{t}$, and selects an action $a_{t}$ under policy $\pi_{\theta}$.
+This process continues until task completion or step limits are reached, with binary rewards indicating success or failure.
+4.2 Framework Overview
+Web-CogReasoner is a knowledge-driven reasoning system built on LVMs and trained on Web-CogDataset (Section 3.3).
+As illustrated in Figure 3, it tackles complex web tasks by generating a Knowledge-driven Chain-of-Thought (CoT), in which each reasoning stage is explicitly grounded in a layer of Web-CogKnowledge.
+Figure 3: Illustrating the Knowledge-driven Chain of Thought (CoT) process.
+Knowledge-driven CoT Reasoning
+The core of Web-CogReasoner is a structured chain-of-thought (CoT) reasoning process (Figure 3), decomposed into three layers: Factual (identifying page elements and states: “What is on the page?”), Conceptual (inferring roles and interactions: “What does it mean?”), and Procedural (planning goal-directed steps: “How to accomplish the task?”).
+This layered reasoning maps task prompts to executable actions: $\textbf{Task Prompt}\rightarrow\textbf{Knowledge-driven CoT}\rightarrow\textbf{Plan}\rightarrow\textbf{Action}$.
+The agent initiates with a Task and observes the webpage (Observation Space).
+This structured reasoning guides the Planning phase, which decomposes the task and formulates a step-by-step strategy.
+The process culminates in a concrete Action to be executed on the webpage.
+5 Experiment
+5.1 Experimental Setup
+Models and Baselines
+Our proposed Web-CogReasoner builds on Qwen2.5-VL-7B (Bai et al., 2025), extended with a knowledge-induced reasoning framework.
+Training details are provided in Appendix A.3.1.
+For comparison, we include diverse baselines: the vanilla Qwen2.5-VL-7B (zero/few-shot), OpenWebVoyager (He et al., 2024b) in IL and Max variants, the open-source UI-TARS-7B-SFT (Qin et al., 2025), and two powerful commercial LVMs, Claude Sonnet 4 and Gemini 2.5 Pro.
+Together these cover foundational LLMs, end-to-end web agents, and general-purpose multimodal models.
+Datasets
+Web-CogReasoner is trained via supervised fine-tuning on the curated Web-CogDataset (Sec.
+3.3), which aligns samples with Factual, Conceptual, and Procedural knowledge using screenshots, accessibility trees, and reasoning trajectories.
+Evaluation is conducted on four datasets.
+Our custom Web-CogBench (Sec.
+3.4) assesses cognitive dimensions of Memorizing, Understanding, and Exploring.
+VisualWebBench (Liu et al., 2024b) provides 1.5K curated tasks across 139 real websites, testing grounding and reasoning in diverse settings.
+WebVoyager (He et al., 2024b) contains 643 queries across 15 seen sites, measuring performance in familiar environments.
+Online Multimodal-Mind2Web (Deng et al., 2023) evaluates cross-task and cross-website generalization with queries from both known and unseen domains.
+5.2 Main Results
+Table 2: Performance evaluation on the Web-CogBench benchmark.
+Model Element Attribute Rec Next Page Pre Source Element Pre Element Understanding Claude Sonnet 4 79.7 93.5 62.5 62.8 Gemini 2.5 Pro 79.8 94.6 84.4 62.6 Qwen2.5-VL-7B 53.2 83.9 65.6 60.0 UI-TARs-7B-SFT 63.5 88.0 31.3 48.0 Web-CogReasoner (Ours) 91.4 93.5 87.5 69.2
+WebPage Understanding User Intent Pre Popup Close Single Step Exp Overall
+Claude Sonnet 4 54.3 64.7 100 96.8 76.8
+Gemini 2.5 Pro 73.5 51.9 96.6 98.4 80.4
+Qwen2.5-VL-7B 62.0 51.9 91.4 90.3 69.8
+UI-TARs-7B-SFT 48.0 32.4 25.9 33.9 46.4
+Web-CogReasoner (Ours) 79.0 61.4 98.3 95.2 84.4
+Results on Web-CogBench
+We evaluate Web-CogReasoner on Web-CogBench, which assesses foundational knowledge and cognitive reasoning across twelve web tasks.
+As Table 2 shows, our model outperforms both commercial and open-source baselines, owing to the integration of structured Web-CogKnowledge (factual, conceptual, procedural) with Knowledge-driven Chain-of-Thought reasoning.
+This combination enables accurate perception of web elements and informed, step-wise decision-making.
+The strong synergy between visual recognition and cognitive reasoning allows Web-CogReasoner to excel on diverse tasks and generalize effectively to web navigation scenarios.
+We also provide a detailed LVM evaluation protocol and inter-rater reliability analysis in Appendix A.4 to validate the robustness of our scoring metrics.
+Table 3: Performance evaluation on the VisualWebBench benchmark.
+Model Perception-Oriented Tasks Perception Avg Reasoning-Oriented Tasks Reasoning Avg Overall Avg
+WebQA HeadOCR OCR Element Ground Action Prediction Action Ground
+Claude Sonnet 4 73.3 72.6 96.2 80.7 81.1 96.1 96.3 91.2 85.9
+Gemini 2.5 Pro 74.9 70.8 95.1 80.3 91.8 96.8 90.3 93.0 86.6
+Qwen2.5-VL-7B 70.8 71.7 81.4 74.6 77.5 86.8 68.0 77.4 76.0
+UI-TARs-7B-SFT 71.3 78.7 97.2 82.4 91.8 91.8 85.4 89.7 86.0
+Web-CogReasoner (Ours) 67.2 72.6 97.0 79.0 96.4 96.1 88.4 93.6 86.3
+Results on VisualWebBench
+We evaluate visual understanding on VisualWebBench (Liu et al., 2024b).
+As Table 3 shows, Web-CogReasoner achieves the highest average score (86.3%), slightly above UI-TARs (86.0%).
+However, UI-TARs performs poorly on Web-CogBench (46.4%), highlighting that strong visual perception alone does not ensure robust cognitive reasoning.
+In contrast, Web-CogReasoner consistently excels across both visual and reasoning benchmarks, demonstrating the effective integration of precise visual perception with structured knowledge-driven reasoning — a dual capability essential for reliable web agents.
+Table 4: Task success rates on the WebVoyager. The ”Overall” score is the average success rate.
+Agent Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
+Claude Sonnet 4 26.7% 87.8% 48.8% 69.8% 68.3% 2.3% 45.5% 83.3%
+Gemini 2.5 Pro 60.0% 63.4% 62.8% 67.4% 68.3% 9.1% 56.8% 73.8%
+Qwen2.5-VL-7B 0.0% 0.0% 0.0% 4.7% 0.0% 0.0% 0.0% 2.3%
 […]
-$(h_{t},a_{t})=\pi_{\theta}(\cdot|K,I,Q,o_{1},h_{1},a_{1},\dots,o_{t})$
-Finally, the overall performance is assessed using a binary reward function $R$, which returns 1 only if the agent successfully completes the user-specified task at the final step, and 0 otherwise.
-4.2 Web-CogReasoner Framework
-We introduce Web-CogReasoner, a knowledge-driven reasoning system built on LVMs and trained on the Web-CogDataset in Section3.3.
-It solves complex web tasks by performing chain-of-thought reasoning grounded in knowledge acquired during training.
-We define web agent actions set based on common user behaviors observed during web browsing.
-At each interaction step, the model receives a screenshot of the current webpage along with its corresponding AX Tree.
-The model integrates visual, structural representations with other contextual cues and history trajectories to predict the next action.
-Once an action is predicted, it is executed by an automated web agent.
-This process iterates until the task is completed or reach the maximum step limit.
-The following section details the design of our agent and reasoning framework:
-Observation Space
-To effectively understand and interact with complex web interfaces, our web agent takes as input both webpage screenshots and corresponding AX Tree.
-The screenshot provides rich visual context essential for grounding visual reasoning, while the AX Tree offers a structured semantic representation of page elements.
-Specifically, the AX Tree encodes key properties such as role (indicating the type of an element), name (how the element should be referred to), description (additional descriptive content), and state (e.g., whether a checkbox is checked or a dropdown is expanded).
-Action Space
-To enable web agents to operate in real-world environments, we define an action space (Table 3) grounded in common user interactions during web browsing.
-Each action $a_{t}$ operates on a target element identified by a unique ID from the current observation $x_{t}$.
-For agents using AX Tree representations, this ID refers to a node within the tree.
+[…]
+Web-CogReasoner (Ours) 26.7% 31.7% 32.6% 34.9% 29.3% 2.3% 15.9% 54.8%
+BBC News Cambridge Dictionary Google Flights Google Map Huggingface Wolfram Alpha Overall
+Claude Sonnet 4 23.8% 37.2% 4.8% 80.5% 48.8% 82.6% 47.7%
+Gemini 2.5 Pro 52.3% 76.7% 4.8% 75.6% 58.1% 82.6% 54.9%
+Qwen2.5-VL-7B 0.0% 11.6% 0.0% 2.4% 7.0% 2.2% 2.2%
+[…]
+[…]
+Web-CogReasoner (Ours) 14.3% 55.8% 9.5% 39.0% 37.2% 39.1% 30.2%
+Table 5: Performance comparison on the Online-Mind2Web under cross-task and cross-websites
+Agent Cross-task (Unseen Task) Cross-web (Unseen Websites)
+Entertainment Shopping Travel Overall Entertainment Shopping Travel Overall
+Claude Sonnet 4 44.9% 35.3% 40.0% 40.2% 45.5% 6.7% 14.0% 21.7%
+Gemini 2.5 Pro 46.9% 35.3% 28.3% 37.5% 42.4% 10.0% 23.3% 25.5%
+OpenWebVoyagerMax 22.4% 29.4% 15.2% 20.5% 3.0% 8.7% 23.3% 11.7%
+Qwen2.5-VL-7B 2.2% 0.0% 0.0% 1.0% 3.0% 0.0% 0.0% 1.0%
+OpenWebVoyagerIL 8.2% 5.9% 4.3% 6.3% 3.0% 5.8% 4.7% 6.6%
+Web-CogReasoner (Ours) 16.3% 23.5% 15.2% 17.0% 12.1% 7.7% 9.3% 10.1%
+Results on Online Web Tasks
+[…]
+Table 6: Average steps per successful task across different benchmarks.
+Agent WebVoyager Mind2Web Cross-Task Mind2Web Cross-Web Final Avg
+Claude 7.35 10.89 11.04 9.76
+Gemini 6.68 7.74 10.30 8.24
+[…]
+Qwen2.5-VL-7B 7.69 12.00 13.00 10.9
+[…]
+Ours 4.73 7.37 8.89 7.00
+Results on Average Steps
+Table 6 reports the average number of steps for successful online tasks.
+Our approach consistently achieves high efficiency, particularly in cross-domain scenarios, indicating that the model effectively balances streamlined task execution with robust generalization to unseen environments.
+5.3 Ablation Study
+To empirically validate the effectiveness of our Web-CogKnowledge Framework, we conduct a two-fold ablation study.
+First, we evaluate the cumulative gains of our curriculum learning strategy (Table 7).
+Second, to address specific inquiries regarding the necessity of each knowledge layer and the reasoning mechanism, we provide a detailed component analysis on both Web-CogBench and WebVoyager (Tables 8 and 9).
+Table 7: Cumulative Gains: Impact of progressive knowledge training on Web-CogBench.
+Model Configuration Memorizing Understanding Exploring Overall
+Qwen2.5-VL-7B (Base Model) 67.6 61.0 77.9 69.8
++ Factual Knowledge (S1) 85.5 (+17.9) 64.2 60.1 72.1
++ Conceptual Knowledge (S2) 88.1 75.5 (+11.3) 65.8 78.3
++ Procedural Knowledge (S3) 90.8 74.1 85.0 (+19.2) 84.4
+Cumulative Impact of Curriculum Learning
+We evaluate the roles of Factual, Conceptual, and Procedural knowledge through an ablation study that incrementally augments the base Qwen2.5-VL-7B model and measures performance on Web-CogBench.
+Factual Knowledge provides the perceptual grounding needed for accurate element recognition, strengthening the Memorizing dimension.
+Conceptual Knowledge introduces semantic structure and functional understanding, improving the Understanding dimension and enabling basic multi-step behaviors.
+Procedural Knowledge adds goal-directed planning and execution, yielding major gains in the Exploring dimension.
+Qualitative examples illustrating how perception, interpretation, and planning evolve across stages are presented in Appendix A.5.
+Together, the results show that each knowledge layer is essential, and their integration is key to Web-CogReasoner’s cognitive robustness and performance.
+Table 8: Component Analysis on Web-CogBench: Validating hierarchical dependency.
+Model Memorizing Understanding Exploring Overall
+Base model 67.6 61.0 77.9 69.81
+S1 only 85.5 64.2 60.1 72.12
+S2 only 59.88 68.03 60.00 61.96
+S3 only 52.82 46.40 78.00 60.66
+S1+S2 88.1 75.5 65.8 78.33
+S1+S3 85.11 53.53 82.31 76.17
+S2+S3 64.87 69.74 81.41 72.29
+S1+S2+S3 (Full) 90.8 74.1 85.0 84.44
+Table 9: Impact of Knowledge & KCoT on WebVoyager, real-world online tasks.
+Model Amazon Cambridge Coursera GitHub Overall
+S1 only 12.19% 25.58% 14.28% 7.14% 12.67%
+S3 only 17.07% 11.62% 16.66% 14.28% 13.14%
+S1+S3 29.26% 34.88% 28.57% 16.66% 23.47%
+S1+S2+S3 (w/o KCoT) 19.51% 51.16% 26.19% 23.80% 25.35%
+S1+S2+S3 (w/ KCoT) 31.7% 55.8% 54.8% 29.3% 42.9%
+Hierarchical Dependency of Knowledge
+To demonstrate that our stages are not merely isolated skills but hierarchically dependent layers, we present a detailed breakdown in Table 8 and Table 9.
+  • Low-level Knowledge is a Prerequisite: As seen in Table 8, single-stage models (S2 only, S3 only) perform poorly on comprehensive metrics.
+Crucially, adding Factual training (S1) significantly boosts the performance of higher-level stages.
+For example, on WebVoyager (Table 9), combining S1 with S3 nearly doubles the success rate compared to S3 alone (23.47% vs.
+13.14%).
+This proves that procedural exploration (S3) cannot function effectively without accurate factual grounding (S1).
+  • Integration is Critical: While specific stages excel at their corresponding metrics (e.g., S3 on Exploring), only the fully integrated model (S1+S2+S3) achieves robust performance across all dimensions, confirming that complex web agents require a complete cognitive stack.
+Reasoning Activation via KCoT
+Finally, we investigate the role of our reasoning framework.
+While the full combination of data (S1+S2+S3) builds a strong latent representation, explicit reasoning is required to utilize it.
+As shown in Table 9, removing the Knowledge-driven Chain-of-Thought (w/o KCoT) causes a sharp drop in online success rate from 42.9% to 25.35%.
+This indicates that KCoT acts as a crucial activator, bridging the gap between possessing knowledge and applying it dynamically for decision-making.
+6 Conclusion
+We present Web-CogReasoner, a cognitive-inspired framework for web agents that systematically instills Factual, Conceptual, and Procedural Knowledge, following Bloom’s Taxonomy.
+By coupling the Web-CogKnowledge Framework with the Web-CogDataset and Web-CogBench, our approach enables interpretable, step-wise reasoning through knowledge-driven Chain-of-Thoughts, yielding strong performance on complex web navigation and instruction-following tasks.
+Ablation and qualitative analyses confirm the indispensability of each knowledge stage, demonstrating how a structured curriculum produces robust perceptual and cognitive capabilities.
+While current results rely on imitation learning, future work aims to integrate reinforcement learning to enhance exploration, generalization, and autonomous discovery of procedural knowledge, advancing toward truly adaptive and self-directed web agents.
+Appendix A Appendix
+A.1 Web Cognition and Data Details
+A.1.1 Bloom’s Taxonomy
+  • 1.
+    Factual Knowledge: The foundational layer, encompassing the basic, discrete elements of a discipline that a student must know, such as essential terminology and specific, isolated details.
+  • 2.
+    Conceptual Knowledge: The synthesis of factual elements into a coherent, organized structure.
+This level focuses on the interrelationships between basic elements, including knowledge of classifications, principles, generalizations, theories, and models.
+  • 3.
+    Procedural Knowledge: The knowledge of how to perform a task or inquiry.
+This involves an understanding of specific skills, algorithms, techniques, and methods, representing a shift from ”knowing-what” to ”knowing-how.”
 Table 10: Action space of Web-CogReasoner for web interaction.
@@
 Wait Wait for one second before proceeding
-Knowledge-driven CoT Reasoning
-Figure 4: An overview of the Web-CogReasoner framework, illustrating the Knowledge-driven Chain of Thought (CoT) process.
-The agent initiates with a Task and observes the webpage (Observe Space).
-The Knowledge-driven COT Reasoning module then systematically leverages different types of Web-CogKnowledge—Factual Knowledge (blue), Conceptual Knowledge (green), and Procedural Knowledge (yellow).
-This structured reasoning guides the Planning phase, which decomposes the task and formulates a step-by-step strategy.
-The process culminates in a concrete Action to be executed on the webpage.
-To instantiate the hierarchical Web-CogKnowledge Framework in a reasoning-capable agent, we introduce the Web-CogReasoner, a structured cognitive web-agent that generates a Knowledge-driven Chain-of-Thought (CoT) to guide decision-making in web environments.
-As illustrated in Figure 4, this CoT is not a monolithic or opaque text block; rather, it is a scaffolded reasoning template, where each segment is explicitly grounded in a distinct type of web knowledge.
-This design ensures interpretability, traceability, and faithful knowledge alignment throughout the reasoning process.
-The CoT construction follows a layered progression, with each stage corresponding to a cognitive function defined in the Web-CogKnowledge taxonomy:
-  • Factual Knowledge (highlighted in blue) constitutes the foundation of reasoning by anchoring the model in observable page facts.
-This phase populates the Webpage Layout Description and Key Element Analysis sections.
-The agent extracts concrete interface features such as DOM structure, ARIA attributes (e.g., role=‘search’), accessible names, and current page states.
-At this stage, the agent is essentially answering the question: “What is on the page?”
-  • Conceptual Knowledge (highlighted in green) builds upon the factual layer by introducing semantic understanding.
-Within the same sections, the model interprets the purpose of key UI components and predicts their behavior under user interaction.
-For instance, it explains the functional intent of the search bar (“allows users to input a location”) and anticipates changes to the page layout (“the list of stores will update after input”).
-This phase reflects the agent’s capacity to comprehend: “What does this element mean?” and “How does it contribute to the task?”
-  • Procedural Knowledge (highlighted in yellow) informs goal-oriented planning and task decomposition.
-It is primarily instantiated in the Task Recap, Task Decomposition, Step-by-Step Reasoning, and Final Action Summary sections.
-The Reasoner infers the user’s intent from the task prompt, predicts the next sub-goal for interaction (e.g., “Users will type in their desired location”), and generates a sequential plan:
-    1. 1.
-      Enter the ZIP code 90028 into the search bar;
-    2. 2.
-      Review the updated list of Apple Stores;
-    3. 3.
-      (Fallback) Click the “Complete store list” if results are missing.
-    This layer translates semantic understanding into executable strategies—effectively answering “How should the task be accomplished?”
-By modularizing the CoT along these knowledge axes, Web-CogReasoner achieves structured planning grounded in a cognitively coherent framework.
-Unlike conventional end-to-end models that produce opaque or hallucinated justifications, each reasoning step here is explicitly traceable to factual evidence, semantic inference, or procedural logic.
-This alignment reduces ungrounded output and enhances interpretability.
-As shown in Figure 4, the color-coded CoT segments (yellow, purple, blue) illustrate how the agent transitions from perception to comprehension to action. This workflow encapsulates the chain:
-[…]
-Through this process, the Web-CogReasoner becomes a practical embodiment of Web-CogKnowledge, leveraging its hierarchical structure to support robust, transparent, and goal-directed web reasoning.
-4.3 Training
-We employ a multi-stage Imitation Learning strategy to train our model on the Web-CogDataset, utilizing Qwen2.5-VL-7B bai2025qwen2 as the base model.
-Each stage is aligned with a distinct layer of the Web-CogKnowledge Framework: (1) the first knowledge content learning focuses on acquiring Factual Knowledge and Conceptual Knowledge, enabling the model to interpret web content and semantics; (2) the second cognitive process emphasizes Procedural Knowledge, training the model to plan and execute multi-step web interactions.
-To accommodate the increased complexity of the final stage, which involves multi-image inputs and extended reasoning, we configure training with a maximum sequence length of 8K and a batch size of 1 with gradient accumulation of 16 steps.
-All experiments are conducted on a cluster of 8 x NVIDIA A800 80GB GPUs.
-5 Experiment
-5.1 Experimental Setup
-5.1.1 Models and Baselines
-To comprehensively evaluate Web-CogReasoner, we compare its performance against several representative and state-of-the-art baselines.
-Our proposed model, Web-CogReasoner, is built upon the Qwen2.5-VL-7Bbai2025qwen2 large multimodal model (LMM) as its foundational architecture.
-As detailed in Section 4, Web-CogReasoner extends Qwen2.5-VL by integrating a knowledge-induced cognitive reasoning framework, which leverages a structured knowledge base (described in Section 3) to guide chain-of-thought (CoT) reasoning for more robust and generalizable web navigation.
-Specifically, we performed supervised fine-tuning on Qwen2.5-VL using our curated datasets to imbue it with the proposed knowledge-driven capabilities.
-[…]
-These baselines collectively represent a diverse spectrum of current web agent methodologies, including foundational LLMs, interactive LLM-based agents, and powerful general-purpose LLM/LMMs leveraging prompting strategies.
-This allows for a thorough and multifaceted evaluation of Web-CogReasoner’s effectiveness and its unique advantages conferred by knowledge-induced cognitive reasoning.
-5.1.2 Datasets
-Training Dataset
-In the training phase, we utilized a comprehensive cognitive dataset——Web-CogDataset,which is meticulously constructed as detailed in Section 3.3.
-This dataset is specifically designed to imbue Web-CogReasoner with knowledge-induced cognitive reasoning capabilities.
-It encompasses a rich variety of samples aligned with Factual, Conceptual, and Procedural knowledge dimensions, including diverse web page screenshots, corresponding accessibility trees, and detailed interaction trajectories with associated reasoning chains.
-Its comprehensive and structured nature is crucial for enabling the model to learn and perform sophisticated web navigation tasks.
-Evaluation Datasets
-To rigorously evaluate Web-CogReasoner’s performance and generalization capabilities, we employ a multi-faceted evaluation strategy utilizing various prominent datasets.
-We utilize the following datasets for evaluation:
-  • Web-CogBench: As elaborated in Section 3.4, this custom-built benchmark is meticulously designed to assess distinct knowledge-driven cognitive abilities: Memorizing, Understanding, and Executing.
-Aallowing for granular evaluation of how our knowledge-induced reasoning framework contributes to these specific competencies.
-  • VisualWebBench liu2024visualwebbench : A human-curated multimodal benchmark covering 1.5K examples across 139 real websites and 87 sub-domains, designed to evaluate MLLMs’s grounding and reasoning abilities in complex, web-based scenarios.
-  • WebVoyager he2024openwebvoyager : This dataset comprises tasks on 15 websites that were seen during training and includes a total of 643 online web task queries.
-It primarily assesses the agent’s ability to perform tasks on familiar web environments.
-  • Online Multimodal-Mind2Web deng2023mind2web : This includes two complementary subsets.
-The Cross-Task Test Set contains 112 queries from 33 websites seen during training, evaluating generalization to new tasks within known structures.
-The Cross-Website Test Set comprises 106 queries from 6 unseen websites across familiar domains (Entertainment, Shopping, and Travel), assessing the agent’s ability to transfer domain knowledge to novel layouts.
-5.2 Evaluation Metrics
-To provide a comprehensive and quantitative assessment of our model’s performance, we employ a set of specific metrics tailored to the nature of our evaluation datasets.
-Web-CogBench Metrics
-For our Web-CogBench, which evaluates distinct cognitive abilities, we use a variety of metrics depending on the task format:
-  • Accuracy: For tasks with a definitive correct answer, such as multiple-choice questions (Next Page Prediction) or single-element identification (Source Element Prediction, Popup Close, Single Step Exploration), we calculate the standard accuracy.
-An answer is considered correct only if it exactly matches the ground truth.
-  • ROUGE-L: For generative tasks that produce open-ended text, such as Element Attribute Recognition, we use ROUGE-L to measure the F1-score of the longest common subsequence between the predicted text and the ground truth.
-This metric effectively evaluates content overlap for short-form text generation.
-  • LVM Judge: For tasks requiring qualitative assessment of generated content, such as Element Understanding and WebPage Understanding, a simple string match is insufficient.
-Therefore, we employ a powerful Large Vision Model (LVM) as an automated judge.
-The LVM is prompted to evaluate the quality, relevance, and accuracy of the model’s generated description against the ground truth, returning a numerical score on a 5-point scale.
-This approach provides a scalable and consistent method for evaluating complex, open-ended visual reasoning.
-Online Web Tasks Metric
-For online evaluation on benchmarks like WebVoyager and Multimodal-Mind2Web, the primary metric is the Task Success Rate (SR).
-A task is considered a success if and only if the agent successfully completes all required sub-steps and achieves the final goal specified in the instruction within the maximum allowed number of steps.
-The final success rate is the percentage of successfully completed tasks over the total number of tasks in the test set.
-This metric provides a clear and direct measure of the agent’s practical utility and effectiveness in real-world, dynamic environments.
-5.3 Main Result
-To validate the effectiveness of our proposed Web-CogReasoner and its Web-CogKnowledge Framework inspired by Bloom’s taxonomy, we conduct a comprehensive set of experiments across several challenging benchmarks.
-We evaluate our model’s performance in three key areas: fine-grained cognitive capabilities, visual understanding, and practical generalization on live online tasks.
-Our model is compared against leading proprietary agents, such as Claude Sonnet 4 and Gemini 2.5 Pro, as well as state-of-the-art open-source models.
-Table 4: Performance evaluation on the Web-CogBench benchmark.
-[…]
-[…]
-[…]
-[…]
-[…]
-Overall
-Claude Sonnet 4 54.3 64.7 100 96.8 76.8
-Gemini 2.5 Pro 73.5 51.9 96.6 98.4 80.2
-Qwen2.5-VL-7B 62.0 51.9 91.4 90.3 69.8
-UI-TARs-7B-SFT 48.0 32.4 25.9 33.9 46.4
-Web-CogReasoner (Ours) 79.0 61.4 98.3 95.2 84.4
-Results on Web-CogBench
-We first evaluate Web-CogReasoner on our proposed Web-CogBench, a benchmark specifically curated to assess the foundational knowledge and cognitive reasoning capabilities of web agents.
-It encompasses twelve diverse tasks ranging from element recognition to user intent inference, in alignment with our two-stage cognitive framework.
-As presented in Table 4, Web-CogReasoner achieves state-of-the-art performance with an overall accuracy of 82.9%, surpassing both commercial and open-source baselines.
-Notably, it outperforms the powerful proprietary model Gemini 2.5 Pro (77.6%) and the strong open-source model Qwen2.5-VL-7B (66.9%), with substantial gains in high-level reasoning tasks such as Source Element Prediction (87.5% vs.
-84.4%) and WebPage Understanding (79.0% vs.
-73.5%).
-These results confirm the effectiveness of our design, which couples structured knowledge acquisition (factual, conceptual, procedural) with Knowledge-driven Chain-of-Thought (CoT) reasoning to form a solid cognitive foundation for agents.
-Importantly, we find that strong performance on Web-CogBench is highly indicative of overall agent effectiveness across web tasks.
-For example, Gemini performs best on Web-CogBench among baseline models and also exhibits the best overall task completion rate in live-agent evaluations, in table 6 and table 7.
-However, we further observe that foundational model capabilities—particularly in visual perception—play a critical role in enabling such high-level cognition.
-Agents with weaker visual recognition skills exhibit reduced performance even when endowed with strong reasoning mechanisms, as reflected in additional evaluation results from VisualWebBench, in table 5.3, especially in perception-oriented tasks.
-In summary, Web-CogBench serves not only as a direct measure of an agent’s reasoning capacity, but also implicitly requires strong perceptual grounding.
-Excelling on Web-CogBench implies that an agent possesses both the knowledge reasoning ability and the necessary perceptual competence—making it a highly reliable proxy for real-world web exploration effectiveness.
-Table 5: Performance evaluation on the VisualWebBench benchmark. Tasks are grouped into Perception-Oriented and Reasoning-Oriented categories.
-[…]
-[…]
-[…]
-[…]
-Ground
-Claude Sonnet 4 73.3 72.6 96.2 80.7 81.1 96.1 96.3 91.2 85.9
-Gemini 2.5 Pro 74.9 70.8 95.1 80.3 91.8 96.8 90.3 92.9 86.6
-Qwen2.5-VL-7B 70.8 71.7 81.4 74.6 77.5 86.8 68.0 77.4 76.0
-UI-TARs-7B-SFT 71.3 78.7 97.2 82.4 91.8 91.8 85.4 89.7 86.0
-Web-CogReasoner (Ours) 67.2 72.6 97.0 79.0 96.4 96.1 88.4 93.6 86.3
-Results on VisualWebBench
-To evaluate the model’s visual understanding capabilities, we conduct experiments on the VisualWebBench benchmark liu2024visualwebbench .
-The results, summarized in Table 5.3, expose the limitations of visually-focused benchmarks and motivate the development of our proposed Web-CogBench.
-Our Web-CogReasoner achieves the highest average score of 86.3%, marginally outperforming UI-TARs (86.0%).
-While this may suggest comparable performance at first glance, a closer examination reveals a stark contrast: UI-TARs performs significantly worse on Web-CogBench, with an overall score of only 48.2% (see Table 4).
-This discrepancy underscores a critical insight — excelling in tasks such as OCR and element localization does not imply competence in cognitive reasoning.
-High scores on visual benchmarks like VisualWebBench can obscure deeper limitations in understanding and decision-making.
-In contrast, Web-CogReasoner demonstrates consistent strength across both visual and reasoning benchmarks, confirming that our framework effectively integrates visual perception with advanced cognitive reasoning.
-This dual capability is essential for building robust and reliable web agents.
-Results on Online Web Tasks
-Beyond offline benchmarks, we evaluated Web-CogReasoner’s practical utility and generalization ability on live, online web tasks using the WebVoyager he2024openwebvoyager and Online Multimodal-Mind2Web deng2023mind2web datasets.
-These experiments test the agent’s ability to apply its knowledge and reasoning to complete complex, multi-step and live tasks on unseen websites and scenarios, which is the ultimate goal for any autonomous web agent.
-Notably, we did not include UI-TARs-7B-SFT qin2025ui in the online evaluation due to the absence of publicly available online inference scripts.
-Therefore, we focus our comparison on the OpenWebVoyager he2024openwebvoyager series, which are specifically designed for web task execution.
-Table 6: Task success rates on the WebVoyager test set. We compare our model, Web-CogReasoner, against several baselines across 14 websites. The "Overall" score is the average success rate.
-Agent Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
-Claude Sonnet 4 26.7% 87.8% 48.8% 69.8% 68.3% 2.3% 45.5% 83.3%
-Gemini 2.5 Pro 60.0% 63.4% 62.8% 67.4% 68.3% 9.1% 56.8% 73.8%
-Qwen2.5-VL-7B 0.0% 0.0% 0.0% 4.7% 0.0% 0.0% 0.0% 2.3%
-[…]
-[…]
-Web-CogReasoner (Ours) 26.7% 31.7% 32.6% 34.9% 29.3% 2.3% 15.9% 54.8%
-[…]
-[…]
-[…]
-[…]
-[…]
-Overall Alpha
-Claude Sonnet 4 23.8% 37.2% 4.8% 80.5% 48.8% 82.6% 47.7%
-Gemini 2.5 Pro 52.3% 76.7% 4.8% 75.6% 58.1% 82.6% 54.9%
-Qwen2.5-VL-7B 0.0% 11.6% 0.0% 2.4% 7.0% 2.2% 2.2%
-[…]
-[…]
-Web-CogReasoner (Ours) 14.3% 55.8% 9.5% 39.0% 37.2% 39.1% 30.2%
-[…]
-Table 7: Performance comparison on the Mind2Web Online dataset under cross-task and cross-websites
-Agent Cross-task (Unseen Task) Cross-web (Unseen Websites)
-Entertainment Shopping Travel Overall Entertainment Shopping Travel Overall
-Claude Sonnet 4 44.9% 35.3% 40.0% 40.2% 45.5% 6.7% 14.0% 21.7%
-Gemini 2.5 Pro 46.9% 35.3% 28.3% 37.5% 42.4% 10.0% 23.3% 25.5%
-OpenWebVoyagerMax 22.4% 29.4% 15.2% 20.5% 3.0% 8.7% 23.3% 11.7%
-Qwen2.5-VL-7B 2.2% 0.0% 0.0% 1.0% 3.0% 0.0% 0.0% 1.0%
-OpenWebVoyagerIL 8.2% 5.9% 4.3% 6.3% 3.0% 5.8% 4.7% 6.6%
-Web-CogReasoner (Ours) 16.3% 23.5% 15.2% 17.0% 12.1% 7.7% 9.3% 10.1%
-The Mind2Web dataset (Table 7) serves as a crucial benchmark for evaluating generalization to unseen tasks (cross-task) and unseen websites (cross-web).
-Our model, Web-CogReasoner, records overall scores of 17.0% and 10.1% in these respective settings.
-[…]
-[…]
-5.4 Ablation Study
-Table 8: Ablation study of our Web-CogReasoner on the Web-CogBench.
-We evaluate the impact of each knowledge stage by progressively training the base model.
-"Base Model" is the original Qwen2.5-VL-7B.
-Each subsequent stage adds a new layer of knowledge training.
-Model Configuration Memorizing Understanding Exploring Overall
-Qwen2.5-VL-7B (Base Model) 67.6 61.0 77.9 69.8
-+ Factual Knowledge (S1) 85.5 (+16.9) 64.2 60.1 72.1
-+ Conceptual Knowledge (S2) 88.1 75.5 (+11.3) 65.8 78.3
-+ Procedural Knowledge (S3) 90.8 74.1 85.0 (+7.1) 84.4
-To empirically validate the effectiveness of our hierarchical, Bloom’s Taxonomy-inspired training methodology, we conduct a thorough ablation study.
-Our goal is to isolate and quantify the contribution of each distinct knowledge layer: Factual, Conceptual, and Procedural.
-We progressively build our Web-CogReasoner agent, starting from the base Qwen2.5-VL-7B model and sequentially adding each knowledge stage.
-The performance at each step is evaluated on our Web-CogBench, as its cognitive dimensions (Memorizing, Understanding, Exploring) directly align with our knowledge framework.
-The results are presented in Table 8.
-Impact of Factual Knowledge
-The first stage of our curriculum introduces Factual Knowledge, teaching the agent the "what" of web elements.
-As shown in Table 8, this initial stage provides a substantial performance boost, raising the overall score from 66.9% to 72.1%.
-Critically, the improvement is most pronounced in the Memorizing dimension, which surges from 68.6% to 85.5%.
-This demonstrates that by learning to perceive and identify concrete, factual information (e.g., element roles and names), the agent builds a solid perceptual foundation, which is the first and essential step towards more complex reasoning.
-Impact of Conceptual Knowledge
-Building upon this factual base, the second stage integrates Conceptual Knowledge, which focuses on the "how things relate." This layer enables the agent to understand the relationships between elements and the underlying semantic structure of a webpage.
-The results show another significant leap in performance, with the overall score climbing to 78.3%.
-The most notable gain occurs in the Understanding dimension, which improves by 11.3%.
-This confirms our hypothesis that once an agent can perceive facts, it must then learn to comprehend their context and purpose to effectively interpret a webpage’s functionality.
-Impact of Procedural Knowledge
-The final stage incorporates Procedural Knowledge, teaching the agent "how to act" to achieve a goal.
-This layer completes the Web-CogReasoner’s cognitive architecture, enabling it to formulate and execute plans.
-This stage delivers the final and most critical improvements for task completion, pushing the overall score to our final state-of-the-art result of 85.0%.
-The most dramatic increase is observed in the Exploring dimension, which skyrockets by 7.1%.
-This finding strongly indicates that goal-oriented, procedural training is paramount for converting passive understanding into active, successful task execution.
-In summary, this staged analysis provides compelling empirical evidence for our core thesis.
-Each knowledge layer, from Factual to Conceptual to Procedural, is an indispensable component that progressively builds the agent’s cognitive capabilities, validating our structured, curriculum-based approach to training highly capable web agents.
-5.5 Qualitative Analysis
-Beyond the quantitative improvements shown in our ablation study, a qualitative analysis of the agent’s behavior at each stage offers deeper insights into how our curriculum shapes its cognitive abilities.
-We examine the agents’ performance on a representative task: "Find and add a laptop under \$1000 to the shopping cart on an e-commerce website."
-Base Model (Qwen2.5-VL-7B)
-Without any specialized training, the base model struggles to formulate a coherent plan.
-Its reasoning is often generic and untethered from the specific UI.
-It might correctly identify a "search bar" but fails to execute a meaningful action, or hallucinates actions that are not possible.
-For instance, its thought process might be: "I should search for a laptop," but its action is an ungrounded ‘click "Categories"‘ because it lacks the procedural knowledge to connect intent to a multi-step sequence of actions.
-Stage 1 Agent (+ Factual Knowledge)
-After training on Factual Knowledge, the agent’s perceptual abilities are significantly enhanced.
-It can now accurately identify and label key elements with their correct attributes.
-Its thought process becomes grounded in the facts of the page: "I see a search bar [ID: 25] with the name ’Search products’.
-I see a button [ID: 28] with the name ’Search’." However, it still struggles with planning.
-It understands "what" is on the page but not "why" or "how" to use it.
-It might correctly type "laptop" into the search bar but then get stuck, not understanding that the next logical step is to click the search button to submit the query.
-Stage 2 Agent (+ Conceptual Knowledge)
-With the addition of Conceptual Knowledge, the agent begins to understand the relationships between elements and their purpose.
-Its reasoning graduates from simple identification to semantic interpretation.
-The thought process now reflects this understanding: "The search bar [ID: 25] is for inputting queries.
-The search button [ID: 28] is functionally linked to it and will trigger the search.
-This group of elements forms a ’search component’." This allows it to reliably complete the search and navigate to the results page.
-However, on the results page, it may still struggle with complex procedural logic, such as applying a price filter.
-Full Model (Web-CogReasoner, + Procedural Knowledge)
-The final agent, equipped with Procedural Knowledge, demonstrates goal-oriented planning and execution.
-It seamlessly translates the high-level task into a concrete action sequence.
-Its thought process is now a strategic plan: "Goal: Add laptop under \$1000 to cart.
-Step 1: Type ’laptop’ into search bar [ID: 25].
-Step 2: Click search button [ID: 28].
-Step 3: On the results page, locate the ’Price Range’ filter.
-Step 4: Input ’1000’ into the ’max price’ field [ID: 57].
-Step 5: Identify a suitable product from the filtered list and click its ’Add to Cart’ button [ID: 83]." This demonstrates a complete cognitive loop from perception and understanding to successful action, validating the necessity of the final procedural training stage.
-Appendix A Appendix
-A.1 Bloom’s Taxonomy and Web Cognitive
-A.1.1 Bloom’s Tax
-  • 1.
-    Factual Knowledge: The foundational layer, encompassing the basic, discrete elements of a discipline that a student must know, such as essential terminology and specific, isolated details.
-  • 2.
-    Conceptual Knowledge: The synthesis of factual elements into a coherent, organized structure.
-This level focuses on the interrelationships between basic elements, including knowledge of classifications, principles, generalizations, theories, and models.
-  • 3.
-    Procedural Knowledge: The knowledge of how to perform a task or inquiry.
-This involves an understanding of specific skills, algorithms, techniques, and methods, representing a shift from "knowing-what" to "knowing-how."
-A.1.2 Web data introduction
+A.1.2 Action Space Overview
+Table 10 summarizes the action space of Web-CogReasoner for web interaction tasks.
+The table categorizes actions into two groups: (1) element-specific operations, such as clicking, typing, double-clicking, and scrolling individual elements; and (2) page-level control actions, including navigation commands, task restart, waiting, and final answer submission.
+This structured action space enables the agent to perform a diverse set of interactions, facilitating comprehensive exploration and manipulation of web pages in a controlled and systematic manner.
+A.1.3 Stages of Human Knowledge and Cognitive Development
+Figure 4: How people handle unfamiliar web pages. People learn factual, conceptual, and procedural knowledge to memorize, understand, and explore the web, ultimately completing specific tasks.
 Cognitive science typically classifies human knowledge into three categories—factual, conceptual, and procedural—which correspond to different stages of cognitive development: perceiving, understanding, and executing.
@@
 An illustrative example is shown in Figure4).
-Figure 5: How people handle unfamiliar web pages. People learn factual, conceptual, and procedural knowledge to memorize, understand, and explore the web, ultimately completing specific tasks.
-A.1.3 Data Sourcing
+A.1.4 Web data introduction
+Figure 5 presents the statistics of the selected websites, grouped by category.
+The visualization highlights the relative proportions of different types of websites included in our study, providing an overview of the dataset composition.
+Such a distribution analysis is crucial for understanding potential biases, ensuring coverage across diverse web domains, and evaluating the generalizability of models trained or tested on these websites.
+Figure 5: Statistics of selected websites by category.
+Data Composition and Balance
+To avoid domain overfitting and ensure both interaction depth and broad generalization, Web-CogDataset employs a strategic hybrid data composition:
+  1. 1.
+    Depth via Self-Collected Data: We selected 14 complex websites for high-depth interaction mining to capture intricate logic often missed by general crawls.
+As shown in Figure 5, we strictly maintained category balance across E-commerce, Finance, Developer Tools, and Social Media within this subset to prevent bias toward any single domain.
+  2. 2.
+    Breadth via Open-Source Augmentation: To address the concern of limited domain diversity (e.g., lack of News, Education, or Forums), we incorporated and enhanced large-scale open-source datasets, including MultiUI (Liu et al., 2024a), Mind2Web (Deng et al., 2023), and OpenWebVoyager (He et al., 2024b).
+Notably, MultiUI is derived from FineWeb (Common Crawl), providing massive coverage of general-purpose webpages.
+This combination ensures our model generalizes to the ”wild” web and is not overfitted to specific interaction styles like financial trading or shopping.
+Annotation Reliability
+We validated our automated annotations via double-blind human verification and cross-model consistency checks (e.g., using GPT-4o). As shown in Table 11, the error rate is minimal.
+Table 11: Reliability Check of Web-CogDataset Annotations.
+Annotation Task Human Verification (Acc) Cross-Model Consistency
+Element Attribute 99.2% 98.5%
+Page Change Pred 97.5% 96.8%
+Sub-element Pred 96.8% 95.4%
+Average 97.8% 96.9%
+A.1.5 Data Sourcing
 To collect comprehensive metadata from web pages, we developed a data collection tool based on Playwright.
@@
 A.2 Task definition
-A.2.1 Factual Web Knowledge
-Element Attribute Rec
+A.2.1 Web-CogDataset
+Table 13: Statistics of Web-CogDataset.
+Knowledge Task Statistics
+Subtotal Total
+Factual Web Knowledge Element Attribute Recognition 37K 81K
+Sub-element Prediction 1K
+Page Change Prediction 24K
+Next Page Prediction 18K
+Source Element Prediction 1K
+Conceptual Web Knowledge Element Understanding 40K 62K
+WebPage Understanding 7K
+Caption & QA 15K
+Procedural Web Knowledge User’s Intention Prediction 2K 27K
+Popup Close 1K
+Single-Step Web Task 17K
+Noisy Multi-Step Web Task 7K
+A.2.2 Web-CogBench
+  • Element Attribute Recognition: Given a screenshot with a highlighted interactive element, the model predicts its semantic role (e.g., button, link) and accessible name (e.g., ”Submit”, ”Search”), relying solely on visual cues.
+  • Next Page Prediction: The model predicts the subsequent page that results from interacting with a specific element on the current page.
+To enhance generalization, we designs two types of tasks: multiple-choice questions and open-ended responses.
+  • Source Element Prediction: Given two screenshots, the current page and the resulting target page, then the model identifies which of the visually marked elements on the current page leads to the target, simulating visual cause-and-effect reasoning.
+  • Element Understanding: For a specific interactive element, the model generates an open-ended paragraph that comprehensively describes the element’s Visible Traits (e.g., text, shape, styling), its On-page Location (e.g., header, sidebar, main content), and its likely User-facing Function (e.g., playing a video, navigating to a new page), relying solely on visual context.
+  • WebPage Understanding: Given a full-page screenshot, the model generates a comprehensive overview describing the webpage’s Layout Organization (e.g., header, event information, seating chart, filter panel), Key Element Analysis (e.g., element attribute, description, function, interaction, expected outcome), and a Summary of the WebPage.
+This enables a thorough understanding of the webpage’s structure and functionality.
+  • User’s Intention Prediction: The model infers high-level user intent from a sequence of webpage screenshots representing an interaction trajectory, requiring visual understanding and temporal reasoning.
+The task is built on the MultiModal-Mind2Web (Deng et al., 2023) dataset, mapping screenshot sequences to natural language instructions.
+  • Popup Close: The model identifies and dismisses popups (e.g., notification modals, login forms) on synthesized webpage screenshots, using a dataset of 51 popup components from JS Design [^3] https://js.design/.
+overlaid on OpenWebVoyager (He et al., 2024b) webpages, with combinatorial augmentation of closing strategies for diverse training.
+  • Single Step Exploration: This task is derived from a multi-step trajectory exploration task and has been decomposed into single-step exploration subtasks.
+For each step, the model receives as input the corresponding accessibility tree (AxTree) and a screenshot of the current webpage.
+Based on these observations, the model performs reasoning to generate the appropriate action along with the target object, effectively simulating a realistic single-step web navigation scenario.
+By breaking down complex multi-step interactions into atomic actions, this setup allows for fine-grained evaluation of the agent’s decision-making capabilities and supports systematic analysis of its performance in web exploration tasks.
+Table 13 summarizes the overall statistics of Web-CogDataset.
+The dataset covers three layers of knowledge—Factual, Conceptual, and Procedural— each corresponding to distinct families of web reasoning tasks.
+Factual Web Knowledge focuses on recognizing attributes, predicting element relationships, and modeling page transitions, totaling 81K instances.
+Conceptual Web Knowledge emphasizes semantic understanding and cross-element comprehension, with 62K instances.
+Procedural Web Knowledge involves action-oriented reasoning tasks, such as predicting user intentions and executing goal-directed interactions, comprising 27K instances.
+Together, these task distributions reflect the hierarchical design of Web-CogDataset and ensure balanced coverage from low-level perception to high-level reasoning.
+A.2.3 Factual Web Knowledge
+Element Attribute Recognition
 We define the Element Attribute Recognition task to assess a model’s capability to infer the interactive semantics of web elements exclusively from visual input.
 Given a full-page screenshot with a specific interactive element marked by a red bounding box, the model is tasked with predicting two key attributes:
-  • the semantic role of the highlighted element (e.g., "button", "link", "checkbox"),
-  • the semantic name, which refers to the element’s accessible textual description (e.g., "Submit", "Search", "Next").
+  • the semantic role of the highlighted element (e.g., ”button”, ”link”, ”checkbox”),
+  • the semantic name, which refers to the element’s accessible textual description (e.g., ”Submit”, ”Search”, ”Next”).
 The ground truth for both attributes is derived from the element-level metadata collected as described in Section A.1.5.
 This task simulates the human cognitive ability to interpret the function of web interface components through visual perception alone, without relying on HTML structure or programmatic representations.
-Sub-elements Pre
+Sub-elements Prediction
 We define the Element Sub-element Prediction task to evaluate a model’s ability to infer the hierarchical structure of web interfaces—specifically, to identify the sub-elements that become visible upon interaction with a given parent element, using only visual information.
@@
 This task simulates the human cognitive process of understanding interactive dependencies in a graphical interface—recognizing not only that a component is clickable, but also predicting its dynamic expansion behavior.
-Page Change Pre
+Page Change Prediction
 We define the Page Change Prediction task to evaluate a model’s capability to infer the visual consequences of interacting with a specific web element, relying solely on visual input.
 In this task, the model is presented with a full-page screenshot in which a target interactive element is highlighted by a red bounding box.
-The model is required to in an open-ended format to predict the visual changes that are likely to occur on the page after the element is clicked.The ground truth for this task is obtained from the generated responses of Qwen-VL 72B, which were produced based on visual metadata, as detailed in SectionA.1.3.
+The model is required to in an open-ended format to predict the visual changes that are likely to occur on the page after the element is clicked.The ground truth for this task is obtained from the generated responses of Qwen-VL-72B, which were produced based on visual metadata, as detailed in SectionA.1.5.
 This task is designed to simulate the human cognitive ability to anticipate the dynamic behavior of web interfaces through perception alone—without access to the underlying source code or prior knowledge of the page logic.
-Next Page Pre
+Next Page Prediction
 We define the Next Page Prediction task to evaluate a model’s ability to forecast navigation outcomes.
@@
 This task assesses the agent’s understanding of functional relationships between interface elements and destination pages.
-Source element Pre
+Source element Prediction
 We define the Source Element Prediction task to assess a model’s ability to identify which element on a webpage leads to a specific target page, using only visual input.
@@
 This task simulates the human ability to reason about visual cause-and-effect relationships in web navigation, without relying on code or prior knowledge of page logic.
-A.2.2 Understanding Web Knowledge
+A.2.4 Understanding Web Knowledge
 Element Understanding
@@
   • Embedded Image QA: Given a question grounded in the content of an embedded image within a webpage screenshot, the model must produce an accurate, context-aware answer using only visual information.
-These questions may refer to image content (e.g., "What brand is shown in the ad?") or its function in the UI.
+These questions may refer to image content (e.g., ”What brand is shown in the ad?”) or its function in the UI.
   • Webpage Captioning: The model is tasked with generating an open-ended description of the webpage’s content, layout, and interactive purpose, treating the entire screenshot as input.
@@
   • Webpage QA: Given a full-page screenshot and a natural language question referring to any aspect of the page (e.g., title, layout, purpose, textual content), the model must generate a grounded and precise answer based on visual and spatial information.
-All four subtasks are derived from the Multi-UI [23] dataset, which provides rich annotations for webpage visual elements and user-facing semantics.
+All four subtasks are derived from the Multi-UI (Liu et al., 2024a) dataset, which provides rich annotations for webpage visual elements and user-facing semantics.
 Together, these subtasks measure a model’s ability to perform grounded visual-language understanding at both local (element-level) and global (page-level) scales.
-A.2.3 Procedual Web Knowledge
-User’s intention Pre
+A.2.5 Procedural Web Knowledge
+User’s intention Prediction
 Built on the MultiModal-Mind2Web dataset—which provides natural language instructions, action trajectories, and aligned web page screenshots—we introduce a novel multi-modal task: inferring the user’s high-level intent from a sequence of visual observations.
@@
 Each screenshot $p_{t}$ corresponds to the visual observation at step $t$ of a successful task execution.
-Formally, the model learns a mapping: $f:\{p_{1},p_{2},……,p_{n}\}\rightarrow Q$ where Q is the natural language instruction.
+Formally, the model learns a mapping: $f:\{p_{1},p_{2},\ldots,p_{n}\}\rightarrow Q$ where $Q$ is the natural language instruction.
   2. 2.
@@
 This diversity ensures comprehensive coverage of real-world popup use cases.
-For background webpages, we utilized the OpenWebVoyager[13] dataset, which contains a large number of authentic webpage screenshots with varied layouts and content, providing a rich foundation for synthesizing realistic popup-injected webpages.
+For background webpages, we utilized the OpenWebVoyager(He et al., 2024b) dataset, which contains a large number of authentic webpage screenshots with varied layouts and content, providing a rich foundation for synthesizing realistic popup-injected webpages.
 To construct the training data for this task, we employed the following procedure:
@@
   2. 2.
-    Generating Popup AX Tree: Each popup image was processed using Qwen-VL 2.5 32B to generate an ARIA-compliant AX Tree.
+    Generating Popup AX Tree: Each popup image was processed using Qwen-VL-2.5-32B to generate an ARIA-compliant AX Tree.
 To simulate diverse structural configurations, we randomly modified the index values of popup AX Tree elements and inserted the popup AX Tree into different locations within the original webpage’s AX Tree, resulting in a combined AX Tree that reflects realistic variations in webpage structure.
   3. 3.
-    Generating Popup Closing Strategies: We then instructed Qwen-VL 2.5 32B to identify all n possible methods for closing the popup, based on the popup image and its corresponding AX Tree.
+    Generating Popup Closing Strategies: We then instructed Qwen-VL-2.5-32B to identify all n possible methods for closing the popup, based on the popup image and its corresponding AX Tree.
 Recognizing that, in practical settings, any correct method is sufficient, we applied combinatorial augmentation to the n methods.
@@
 We define the Single-Step Web Task to evaluate a model’s ability to ground high-level user intentions in visual webpage elements.
-Each task instance includes a full-page screenshot from a real-world webpage, a concise natural language instruction (e.g., "Search for a product", "Log into the system"), and several candidate elements marked by red bounding boxes.
+Each task instance includes a full-page screenshot from a real-world webpage, a concise natural language instruction (e.g., ”Search for a product”, ”Log into the system”), and several candidate elements marked by red bounding boxes.
 The model must identify which element, if clicked, would successfully fulfill the given task.
 This setup simulates perceptual grounding of user intent—matching natural language goals to actionable UI targets based solely on visual cues.
-All samples are directly sourced from the Multi-UI [23] dataset, which provides rich, annotated webpage screenshots paired with task descriptions and labeled ground-truth targets.
+All samples are directly sourced from the Multi-UI (Liu et al., 2024a) dataset, which provides rich, annotated webpage screenshots paired with task descriptions and labeled ground-truth targets.
 No trajectory-level annotation or external instruction rewriting is involved.
@@
 Noisy Multi Step Web Task
-To further enhance the original OpenWebVoyager [13] dataset, we incorporate Knowledge-driven Chain-of-Thought (CoT) Reasoning to improve the model’s stepwise understanding and execution, details see Section  4.2 In addition, to simulate realistic interruptions during multi-step web interactions, we propose the Noisy Multi-Step Web Task by augmenting interaction trajectories from OpenWebVoyager.
+To further enhance the original OpenWebVoyager (He et al., 2024b) dataset, we incorporate Knowledge-driven Chain-of-Thought (CoT) Reasoning to improve the model’s stepwise understanding and execution, details see Section  4.2 In addition, to simulate realistic interruptions during multi-step web interactions, we propose the Noisy Multi-Step Web Task by augmenting interaction trajectories from OpenWebVoyager.
 Specifically, for each sample in our Popup Close dataset, a popup window is injected at a specific step (e.g., step $t$) of an existing task trajectory.
@@
 It also provides a challenging benchmark for evaluating agents’ robustness to UI-level noise and their capacity for error recovery.
+A.3 Train Details
+A.3.1 Training
+We employ a multi-phase Imitation Learning strategy to train our model on Web-CogDataset, utilizing Qwen2.5-VL-7B (Bai et al., 2025) as the base model.
+Each phase is aligned with a distinct layer of the Web-CogKnowledge Framework: (1) the first knowledge content learning focuses on acquiring Factual Knowledge and Conceptual Knowledge, enabling the model to interpret web content and semantics; (2) the second cognitive process emphasizes Procedural Knowledge, training the model to plan and execute multi-step web interactions.
+To accommodate the increased complexity of the final phase, which involves multi-image inputs and extended reasoning, we configure training with a maximum sequence length of 8K and a batch size of 1 with gradient accumulation of 16 steps.
+All experiments are conducted on a cluster of 8 x NVIDIA A800 80GB GPUs.
+A.3.2 Performance Under Different Training Strategies
+We could feed these SFT datasets into models with different training strategies. In this section, We investigate how they influence the model’s performance.
+  1. 1.
+    Curriculum Learning Strategy: We fine-tune the model following a curriculum learning paradigm.
+  2. 2.
+    Mixed Multi-task Learning: We directly mix different tasks and apply SFT.
+Through comparative analysis, we find that: Under the same task scenario, models trained via curriculum learning conduct multiple rounds of exploration continuously and do not cease exploration prematurely.
+In contrast, when retrieving task - related data, models trained via mixed training terminate the search within a limited number of attempts if they fail to find results.
+As shown in Figure 7, when faced with the user’s instruction ”Open the most helpful 5 - star reviews of Alpine Ridge”, models trained via mixed training deviate from the domain of task - specific information retrieval and autonomously generate an instruction to ”switch to forums for information”.
+Based on these observations, we conjecture: Curriculum learning trains models incrementally from simple to complex task, which enables them to develop a fundamental understanding of the problem structure.
+They continue to carry out multiple rounds of exploration to ensure comprehensive comprehension and accurate processing, thus avoiding the premature termination of exploration.
+In contrast, mixed training models may be trained on multiple tasks simultaneously.
+The potential interference among different tasks can make models easily influenced by irrelevant tasks when processing a specific one, thereby undermining their ability to focus on task - specific information retrieval and processing.
+Figure 7: Comparison between mixed and curriculum strategies.
+A.4 Evaluation Details
+A.4.1 LVM Evaluation Protocol
+For open-ended generation tasks in Web-CogBench (i.e., Element Understanding, WebPage Understanding, and User Intent Prediction), we utilize a high-capability LVM (GPT-4o) as an automated evaluator.
+The evaluation process involves a strict comparison between the Candidate Model’s Answer and the verified Ground Truth based on image provided in the dataset.
+To ensure fine-grained assessment, we decompose the evaluation into specific cognitive dimensions rather than relying on a single holistic score:
+  • Element Understanding: Assessed on Appearance (visual fidelity), Position (structural context), and Function (interaction purpose).
+  • WebPage Understanding: Assessed on Structure & Layout, Key Element Analysis, and Summary Coherence.
+  • User Intent Prediction: Assessed on Evidence Alignment (visual cue detection), Intent Accuracy, and Reasoning Quality.
+We employ a rigorous 1-5 Likert Scale and normalized to 0–100 for scoring.
+A score of 5 denotes that the candidate ”fully and accurately captures all relevant information present in the Ground Truth,” while lower scores reflect varying degrees of omissions or inaccuracies.
+The evaluator outputs a structured JSON object containing both integer scores and text justifications for each dimension, ensuring the traceability of the results.
+For instance, the Element Understanding task is assessed on:
+System Instruction: You are a meticulous and impartial AI evaluator for a web UI understanding benchmark.
+Your task is to assess the quality of a candidate model’s answer by comparing it strictly against a ground truth and provided image reference.
+Your evaluation must be based exclusively on the information provided in the “Ground Truth Answer” and ”Image”.
+Evaluate the candidate answer on three specific aspects: Appearance, Position, and Function.
+[Ground Truth Answer] {ground_truth} [Candidate Model’s Answer] {model_answer} Evaluation Criteria & Scoring: Score 1: Completely incorrect or missing.
+Score 2: Mostly incorrect, with a minor element of truth.
+Score 3: Partially correct, but misses significant details mentioned in the ground truth.
+Score 4: Mostly correct, with only minor inaccuracies or omissions compared to the ground truth.
+Score 5: Fully and accurately captures all relevant information present in the ground truth.
+Your response MUST be a single, valid JSON object, adhering to the following structure.
+Do not add any text before or after the JSON object.
+{ ”appearance_score”: ¡integer_score¿, ”appearance_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”position_score”: ¡integer_score¿, ”position_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”function_score”: ¡integer_score¿, ”function_justification”: ”¡Your brief justification… referencing the ground truth¿”, ”overall_score”: ¡A final holistic integer score from 1 to 5¿, ”overall_justification”: ”¡A final summary of the model’s performance¿” }
+A.4.2 LVM Judge Reliability
+To mitigate potential biases from a single evaluator, we employed multiple distinct LVMs (including GPT-4o, Claude Sonnet 4, and Gemini 2.5 Pro) to conduct a rigorous inter-rater reliability analysis.
+We calculate the ”Within-1-Point Agreement”, defined as the percentage of instances where scores assigned by different LVM judges differ by no more than 1 point.
+As shown in Table 14, the high agreement rates across different models confirm that our evaluation criteria are robust and model-agnostic.
+Furthermore, the strong correlation with Human Proxy Analysis suggests that our Ground-Truth Anchored protocol effectively aligns automated judgment with human evaluation standards.
+Table 14: Inter-Rater Reliability Analysis of LVM Judge.
+Task Within-1-Point Agreement Human Proxy Analysis
+Element Understanding 98.7% 96.7%
+WebPage Understanding 97.0% 95.4%
+User Intent Prediction 96.0% 94.4%
+A.5 Qualitative Analysis
+To provide deeper insights into Web-CogReasoner’s capabilities, we present a two-part qualitative analysis.
+First, we examine the evolution of cognitive abilities across training stages to validate our curriculum.
+Second, we present a comparative case study on a complex real-world task to demonstrate how our model overcomes knowledge blind spots that trap baseline models.
+A.5.1 Evolution of cognitive abilities
+Beyond the quantitative improvements shown in our ablation study, a qualitative analysis of the agent’s behavior at each stage offers deeper insights into how our curriculum shapes its cognitive abilities.
+We examine the agents’ performance on a representative task: ”Find and add a laptop under \$1000 to the shopping cart on an e-commerce website.”
+Base Model (Qwen2.5-VL-7B)
+Without any specialized training, the base model struggles to formulate a coherent plan.
+Its reasoning is often generic and untethered from the specific UI.
+It might correctly identify a ”search bar” but fails to execute a meaningful action, or hallucinates actions that are not possible.
+For instance, its thought process might be: ”I should search for a laptop,” but its action is an ungrounded ‘click ”Categories” because it lacks the procedural knowledge to connect intent to a multi-step sequence of actions.
+Stage 1 Agent (+ Factual Knowledge)
+After training on Factual Knowledge, the agent’s perceptual abilities are significantly enhanced.
+It can now accurately identify and label key elements with their correct attributes.
+Its thought process becomes grounded in the facts of the page: ”I see a search bar [ID: 25] with the name ’Search products’.
+I see a button [ID: 28] with the name ’Search’.” However, it still struggles with planning.
+It understands ”what” is on the page but not ”why” or ”how” to use it.
+It might correctly type ”laptop” into the search bar but then get stuck, not understanding that the next logical step is to click the search button to submit the query.
+Stage 2 Agent (+ Conceptual Knowledge)
+With the addition of Conceptual Knowledge, the agent begins to understand the relationships between elements and their purpose.
+Its reasoning graduates from simple identification to semantic interpretation.
+The thought process now reflects this understanding: ”The search bar [ID: 25] is for inputting queries.
+The search button [ID: 28] is functionally linked to it and will trigger the search.
+This group of elements forms a ’search component’.” This allows it to reliably complete the search and navigate to the result page.
+However, on the result page, it may still struggle with complex procedural logic, such as applying a price filter.
+Full Model (Web-CogReasoner + Procedural Knowledge)
+The final agent, equipped with Procedural Knowledge, demonstrates goal-oriented planning and execution.
+It seamlessly translates the high-level task into a concrete action sequence.
+Its thought process is now a strategic plan: ”Goal: Add laptop under \$1000 to cart.
+Step 1: Type ’laptop’ into search bar [ID: 25].
+Step 2: Click search button [ID: 28].
+Step 3: On the results page, locate the ’Price Range’ filter.
+Step 4: Input ’1000’ into the ’max price’ field [ID: 57].
+Step 5: Identify a suitable product from the filtered list and click its ’Add to Cart’ button [ID: 83].” This demonstrates a complete cognitive loop from perception and understanding to successful action, validating the necessity of the final procedural training stage.
+A.5.2 Comparative case study
+To validate the necessity of foundational knowledge in handling complex real-world scenarios, we compare the Base Model with Web-CogReasoner on a specific Amazon task.
+Task: ”Find a gaming desktop with Windows 11 Home and 1TB disk.”
+Base Model Failure: The Knowledge Blind Spot
+Lacking explicit knowledge of page layout and element functions, the Base Model literally ”sees” the pixels but ”misses” the affordance.
+  • Observation: The model sees the search results but fails to recognize the sidebar filters as the mechanism to refine the query.
+  • Error: It misinterprets the page state, assuming a re-search is necessary. It enters a logical dead loop of repeatedly clicking the search button.
+  • Action: click [1470] (Search Button) $\to$ Stuck.
+Web-CogReasoner Success: Knowledge-Driven Grounding
+Leveraging learned Factual and Conceptual knowledge, our agent explicitly identifies the functional role of UI components.
+  • Reasoning: The agent identifies the sidebar as a filter section.
+It conceptually maps the user’s ”1TB” requirement to the specific filter element, predicting that clicking it will narrow the results without leaving the page.
+  • Action: click [95] (Filter Link "1 TB") $\to$ Success.
+This comparison highlights that success in complex tasks is not just about planning (Procedural), but requires strictly accurate recognition of element functions (Factual/Conceptual) to ground those plans.